### v3

Abstract
Accurate forecasting of volatility and return quantiles is essential for evaluating financial tail risks such as value-at-risk and expected shortfall. This study proposes an extension of the traditional stochastic volatility model—termed the realized stochastic volatility model—that incorporates realized volatility as an efficient proxy for latent volatility. To better capture the stylized features of financial return distributions, particularly skewness and heavy tails, we introduce three variants of skewed $t$-distributions, two of which incorporate skew-normal components to flexibly model asymmetry. The models are estimated using a Bayesian Markov chain Monte Carlo approach and applied to daily returns and realized volatilities from major U.S. and Japanese stock indices. Empirical results demonstrate that incorporating both realized volatility and flexible return distributions substantially improves the accuracy of volatility and tail risk forecasts.
Keywords: Bayesian estimation, Markov chain Monte Carlo, Realized volatility, Skew-t distribution, Stochastic volatility
1 Introduction
Volatility, defined as the standard deviation or variance of asset returns, evolves stochastically over time and plays a central role in financial risk management. Accurate volatility forecasting is essential for evaluating tail risks such as value-at-risk (VaR) and expected shortfall (ES). Traditionally, two major classes of time-series models have been used for modeling time-varying volatility: the generalized autoregressive conditional heteroskedasticity (GARCH) family (Engle, 1982; Bollerslev, 1986), and the stochastic volatility (SV) model (Taylor, 1986). These models successfully capture key features of financial volatility, including volatility clustering and persistence, and have been extended to account for leverage effects—i.e., the negative correlation between current returns and future volatility. A prominent example is the exponential GARCH (EGARCH) model by Nelson (1991).
More recently, realized volatility (RV), calculated as the sum of squared intraday returns, has gained popularity as a nonparametric and more efficient measure of daily volatility. Comprehensive surveys on RV are provided by Andersen and Benzoni (2009) and McAleer and Medeiros (2008). To model the time-series behavior of RV, researchers have proposed long-memory models such as autoregressive fractionally integrated moving average (ARFIMA) models (Beran, 1994), as well as the heterogeneous autoregressive (HAR) model (Corsi, 2009), which approximates long-memory dynamics using a small number of lags.
Although ARFIMA and HAR models often outperform traditional GARCH and SV models in volatility forecasting, RV estimates are prone to biases caused by market microstructure noise and non-trading hours. Several techniques have been developed to address these issues, including multiscale estimators (Zhang et al., 2005; Zhang, 2006), pre-averaging methods (Jacod et al., 2009), and realized kernel (RK) estimators (Barndorff-Nielsen et al., 2008, 2009). For further refinements and applications, see Aït-sahalia and Mykland (2009), Ubukata and Watanabe (2014), and Liu et al. (2015).
To exploit the advantages of both parametric and nonparametric approaches, hybrid models have been proposed. These include the realized stochastic volatility (RSV) model (Takahashi et al., 2009; Dobrev and Szerszen, 2010; Koopman and Scharth, 2013), as well as the realized GARCH (RGARCH) and realized EGARCH (REGARCH) models (Hansen et al., 2012; Hansen and Huang, 2016). While several studies have compared forecasting performance within each class, comprehensive cross-model evaluations remain limited. Recent studies such as Takahashi et al. (2023, 2024) have attempted to bridge this gap.
Beyond volatility, accurate prediction of financial tail risks requires appropriate modeling of return distributions, which often exhibit skewness and leptokurtosis. While time-varying volatility captures some of this behavior, the conditional distribution of returns may remain non-Gaussian even after accounting for volatility. To address this, skewed Student’s $t$-distributions have been widely applied in volatility models (Abanto-Valle et al., 2015; Kobayashi, 2016). Among them, the generalized hyperbolic (GH) skew-$t$ distribution has shown strong empirical performance in SV (Nakajima and Omori, 2012; Leão et al., 2017) and RSV models (Trojan, 2013; Nugroho and Morimoto, 2014, 2016; Takahashi et al., 2016).
In this paper, we extend the RSV framework by incorporating three types of skew-$t$ distributions: the Azzalini-type skew-$t$ distribution (Azzalini, 1985), the Fernández and Steel (FS) skew-$t$ distribution (Fernández and Steel, 1995), and the GH skew-$t$ distribution (Aas and Haff, 2006). We adopt a Bayesian estimation strategy using Markov chain Monte Carlo (MCMC) simulation and apply the models to daily returns of the Dow Jones Industrial Average (DJIA) and the Nikkei 225 (N225).
Our empirical analysis evaluates the performance of RSV models alongside benchmark SV, EGARCH, and REGARCH models. Volatility forecasts are assessed using the quasi-likelihood loss function, while VaR and ES forecasts are evaluated using the joint loss function proposed by Fissler and Ziegel (2016) and specified in Patton et al. (2019). To test forecast performance, we implement the predictive ability test of Giacomini and White (2006) and the model confidence set procedure by Hansen et al. (2011). The results show that incorporating realized volatility substantially improves forecasting performance. In particular, RSV models with skew-$t$ innovations exhibit superior predictive accuracy, confirming and extending findings from prior research.
The remainder of the paper is organized as follows. Section 2 presents the SV and RSV models, including skewed innovations and a brief overview of realized measures. Section 3 outlines the Bayesian estimation method and forecast evaluation metrics. Section 4 reports the empirical results. Section 5 concludes with a summary of findings and implications.
2 Model
2.1 Stochastic Volatility Model
Let $y_{t}$ denote the daily log return of an asset:
$y_{t}=\log p_{t}-\log p_{t-1},$
where $p_{t}$ is the closing price on day $t$. The return variance is modeled as
$\mathrm{E}[y_{t}^{2}]=\sigma_{t}^{2}=\exp(h_{t}),$
where $h_{t}$ denotes the log volatility, treated as a latent process.
The log volatility $h_{t}$ is assumed to follow a stationary AR(1) process:
$y_{t}$ $=\epsilon_{t}\exp\left(\frac{h_{t}}{2}\right),\quad t=1,\ldots,n,$ (1)
$h_{t+1}$ $=\mu+\phi(h_{t}-\mu)+\eta_{t},\quad t=1,\ldots,n-1,\quad|\phi|<1,$ (2)
$h_{1}$ $\sim\mathcal{N}\left(\mu,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right),$ (3)
$\epsilon_{t}$ $\sim\mathcal{N}(0,1),$ (4)
$\eta_{t}\mid\epsilon_{t}$ $\sim\mathcal{N}\left(\rho\sigma_{\eta}\epsilon_{t},(1-\rho^{2})\sigma_{\eta}^{2}\right).$ (5)
Here, $\phi$ is the persistence parameter, and $\mu$ is the unconditional mean of log volatility. Stationarity is ensured by $|\phi|<1$. The initial state $h_{1}$ is drawn from its unconditional distribution.
The bivariate error term $(\epsilon_{t},\eta_{t})^{\prime}$ follows a jointly normal distribution with correlation $\rho$. This structure allows for a leverage effect: a negative return shock ($\epsilon_{t}<0$) tends to increase future volatility ($h_{t+1}$), reflected by $\rho<0$. This feature captures the empirically observed asymmetry in financial markets, where volatility tends to rise following negative returns.
2.2 Realized Volatility
In the SV model, volatility $\exp(h_{t})$ is treated as a latent variable, since the log-volatility $h_{t}$ is unobserved. With the increasing availability of high-frequency data, RV has emerged as a prominent nonparametric estimator of true volatility.
Let $p(s)$ denote the logarithm of the asset price at time $s$, and suppose it follows the continuous-time diffusion process:
$dp(s)=\mu(s)ds+\sigma(s)dW(s),$ (6)
where $\mu(s)$ and $\sigma(s)^{2}$ are the drift and instantaneous variance, respectively, and $W(s)$ is a standard Brownian motion. The integrated volatility (IV) over a single day is defined as
[…]
Assume that $n$ intraday returns are observed on day $t$, denoted by $\{r_{t},r_{t+1/n},\ldots,r_{t+(n-1)/n}\}$. Then, the realized volatility is defined as
$RV_{t}=\sum_{i=0}^{n-1}r_{t+i/n}^{2}.$ (8)
Under ideal conditions (e.g., no microstructure noise or missing observations), $RV_{t}$ converges to $IV_{t}$ as $n\to\infty$.
However, these ideal conditions are often violated in practice. One source of bias arises from non-trading hours, which are not captured in intraday data. Ignoring these periods leads to an underestimation of daily volatility. To address this, Hansen and Lunde (2005) proposed a scaling adjustment:
$RV_{t}^{HL}=c_{HL}RV_{t},\quad c_{HL}=\frac{\sum_{t=1}^{n}(y_{t}-\bar{y})^{2}}{\sum_{t=1}^{n}RV_{t}},\quad\bar{y}=\frac{1}{n}\sum_{t=1}^{n}y_{t}.$ (9)
This adjustment ensures that the average of the scaled RV matches the sample variance of daily returns.
Another significant bias stems from microstructure noise. Observed prices deviate from the efficient price due to bid–ask bounce, price discreteness, and asynchronous trading, introducing autocorrelation and distortions in high-frequency returns. The bias tends to worsen as the sampling interval becomes shorter. Optimal sampling frequency and alternative estimators have been studied extensively (Aït-Sahalia et al., 2005; Bandi and Russell, 2006, 2008; Liu et al., 2015).
Among various alternatives, the RK estimator (Barndorff-Nielsen et al., 2008) has been widely adopted due to its robustness. It is defined as:
$RK_{t}=\sum_{h=-H}^{H}k\left(\frac{h}{H+1}\right)\gamma_{h},$ (10)
where $H$ is the bandwidth, $k(\cdot)$ is a weight function, and
$\gamma_{h}=\sum_{i=|h|+1}^{m}r_{t-1+i/m}\cdot r_{t-1+(i-|h|)/m}.$ (11)
Barndorff-Nielsen et al. (2009) provide guidelines for choosing the bandwidth $H$.
To account for these biases in modeling, hybrid frameworks such as the RSV and RGARCH models have been developed. In particular, the RSV model of Takahashi et al. (2009) addresses RV bias by jointly estimating model parameters, as detailed in Section 2.3.
2.3 Realized Stochastic Volatility Model
Takahashi et al. (2009) proposed enhancing the SV model by incorporating additional information from RV. In their framework, the daily log-return $y_{t}$ and the logarithm of realized volatility $x_{t}=\log RV_{t}$ are jointly modeled, with the latent log volatility $h_{t}=\log IV_{t}$ serving as a common driver.
The RSV model is specified as:
$x_{t}$ $=\xi+h_{t}+u_{t},\quad t=1,\ldots,n,$ (12)
$y_{t}$ $=\epsilon_{t}\exp\left(\frac{h_{t}}{2}\right),\quad t=1,\ldots,n,$ (13)
$h_{t+1}$ $=\mu+\phi(h_{t}-\mu)+\eta_{t},\quad|\phi|<1,$ (14)
$h_{1}$ $\sim\mathcal{N}\left(\mu,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right),$ (15)
$u_{t}$ $\sim\mathcal{N}(0,\sigma_{u}^{2}),$ (16)
$\epsilon_{t}$ $\sim\mathcal{N}(0,1),$ (17)
$\eta_{t}\mid\epsilon_{t}$ $\sim\mathcal{N}\left(\rho\sigma_{\eta}\epsilon_{t},(1-\rho^{2})\sigma_{\eta}^{2}\right).$ (18)
The error term $u_{t}$ is assumed to be independent of $(\epsilon_{t},\eta_{t})$.
Equation (12) may be generalized as $x_{t}=\xi+\psi h_{t}+u_{t}$, but empirical evidence suggests that estimating $\psi\neq 1$ does not improve forecasting performance. Accordingly, we fix $\psi=1$, following Takahashi et al. (2009).
The parameter $\xi$ adjusts for systematic bias in $\log RV_{t}$. When $\xi=0$, RV is an unbiased estimate of the integrated volatility. In practice, however, $RV_{t}$ may be biased downward due to non-trading hours and upward due to microstructure noise. Hansen and Lunde (2006) show that the net bias can be either positive or negative. Therefore, the sign of $\xi$ reflects the relative magnitude of these two opposing effects.
The distribution of $\epsilon_{t}$ captures residual return behavior after controlling for time-varying volatility. While the SV structure already accounts for volatility clustering and excess kurtosis, it is well documented that financial returns often exhibit additional heavy tails and asymmetry. To model these features more effectively, several studies have incorporated non-Gaussian innovations.
For the SV model, Nakajima and Omori (2012) and Leão et al. (2017) employed the GH skew-$t$ distribution from Aas and Haff (2006). Other approaches include the skew-$t$ distribution by Azzalini and Capitanio (2003) (used in Abanto-Valle et al., 2015), the skew exponential power distribution (Kobayashi, 2016), and the flexible skew-$t$ model by Fernández and Steel (1995), as adopted in Steel (1998). In the context of RSV models, Trojan (2013), Nugroho and Morimoto (2014, 2016), and Takahashi et al. (2016) applied the GH skew-$t$ distribution.
Building on this literature, we extend the RSV model defined in equations (12)–(18)—hereafter referred to as the RSV-N model—by incorporating skewed and heavy-tailed distributions for $\epsilon_{t}$. This extension allows for more accurate modeling of return asymmetry and tail risk, which is essential for reliable risk forecast applications.
2.4 RSV Model with Student’s $t$ Distribution
We extend the RSV-N model by allowing the return innovation $\epsilon_{t}$ to follow a standardized Student’s $t$ distribution. The resulting model, referred to as the RSV-T model, modifies equations (17)–(18) as follows:
$\epsilon_{t}$ $=z_{t}\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}},$ (19)
$z_{t}$ $\sim\mathcal{N}(0,1),$ (20)
$\eta_{t}\mid z_{t}$ $\sim\mathcal{N}\left(\rho\sigma_{\eta}z_{t},(1-\rho^{2})\sigma_{\eta}^{2}\right),$ (21)
where
[…]
Here, $\lambda_{t}$ is an inverse-gamma mixing variable, and $\nu$ is the degrees of freedom parameter controlling tail thickness. As $\nu\to\infty$, the model converges to the Gaussian RSV-N specification.
The use of Student’s $t$ distribution allows the model to accommodate excess kurtosis beyond what is captured by the stochastic volatility component $h_{t}$. For finite $\nu$, the distribution of $\epsilon_{t}$ exhibits heavier tails, improving the model’s ability to reflect extreme return events.
Importantly, the contemporaneous correlation between $\epsilon_{t}$ and $\eta_{t}$ becomes a function of both $\rho$ and $\nu$, given by:
$\mathrm{Corr}[\epsilon_{t},\eta_{t}]=\frac{\mathrm{E}\left[\sqrt{\lambda_{t}}\right]}{\sqrt{\mu_{\lambda}}}\rho,$ (23)
where
$\mathrm{E}\left[\sqrt{\lambda_{t}}\right]=\sqrt{\frac{\nu}{2}}\cdot\frac{\Gamma\left(\frac{\nu-1}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)}.$ (24)
Note that the variance of $\epsilon_{t}$ exists only when $\nu>2$, and its fourth moment exists when $\nu>4$. In empirical applications, $\nu$ is typically estimated from the data.
2.5 RSV Model with GH Skew-$t$ Distribution
Takahashi et al. (2016) extended the RSV-N model in equations (12)–(18) by incorporating the GH skew-$t$ distribution, resulting in the RSV-GH-ST model. This is achieved by replacing $\epsilon_{t}$ and $\eta_{t}$ in (17)–(18) with the following mixture representation:
$\epsilon_{t}$ $=\frac{\beta(\lambda_{t}-\mu_{\lambda})+\sqrt{\lambda_{t}}z_{t}}{\sqrt{\beta^{2}\sigma_{\lambda}^{2}+\mu_{\lambda}}},$ (25)
$z_{t}$ $\sim\mathcal{N}(0,1),$ (26)
$\eta_{t}\mid z_{t}$ $\sim\mathcal{N}\left(\rho\sigma_{\eta}z_{t},(1-\rho^{2})\sigma_{\eta}^{2}\right),$ (27)
where $\lambda_{t}\sim\mathcal{IG}(\nu/2,\nu/2)$, $\mu_{\lambda}=\nu/(\nu-2)$, and
$\sigma_{\lambda}^{2}=\mathrm{Var}[\lambda_{t}]=\frac{2\nu^{2}}{(\nu-2)^{2}(\nu-4)},\quad\nu>4.$ (28)
The denominator in (25) standardizes $\epsilon_{t}$ so that its conditional variance equals $\exp(h_{t})$. When $\beta=0$, the model reduces to the RSV-T specification based on the standard Student’s $t$ distribution.
In this formulation, the skewness and tail thickness of $\epsilon_{t}$ are jointly governed by the parameters $\beta$ and $\nu$. The parameter $\beta$ introduces asymmetry, while $\nu$ controls tail heaviness. As $\nu\to\infty$, or if $\lambda_{t}=1$ for all $t$, the distribution converges to the Gaussian case regardless of $\beta$.
The correlation between $\epsilon_{t}$ and $\eta_{t}$ depends on $\rho$, $\nu$, and $\beta$, and is given by:
$\mathrm{Corr}[\epsilon_{t},\eta_{t}]=\frac{\mathrm{E}[\sqrt{\lambda_{t}}]}{\sqrt{\beta^{2}\sigma_{\lambda}^{2}+\mu_{\lambda}}}\rho,$ (29)
where $\mathrm{E}[\sqrt{\lambda_{t}}]$ is as defined in equation (24).
To illustrate the influence of $\beta$ and $\nu$, Figure 1 plots the simulated densities of $\epsilon_{t}$ under various parameter values. Panel (i) shows that increasing negative values of $\beta$ yield more pronounced left-skewness and heavier tails. In contrast, Panel (ii) indicates that as $\nu$ increases, the distribution becomes more symmetric and approaches normality, exhibiting thinner tails.
Figure 1: Simulated density of the standardized GH skew-$t$ distribution in equation (25). (i) Varying $\beta=0,-1,-2$ with fixed $\nu=10$. (ii) Varying $\nu=24,16,8$ with fixed $\beta=-1$.
2.6 RSV Model with Azzalini Skew-$t$ Distribution
We extend the RSV-N model defined in equations (12)–(18) by adopting the skew-$t$ distribution proposed by Azzalini (1985). [^1] See also Azzalini and Capitanio (2003) and Sahu et al. (2003). The resulting model, referred to as RSV-AZ-ST, modifies the return innovation $\epsilon_{t}$ and its correlation with $\eta_{t}$ as follows:
$\epsilon_{t}$ $=\frac{\delta(z_{0t}-c)+\sqrt{1-\delta^{2}}\,z_{t}}{\sqrt{1-c^{2}\delta^{2}}}\cdot\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}},\quad|\delta|<1,$ (30)
$z_{t}$ $\sim\mathcal{N}(0,1),$ (31)
[…]
$\eta_{t}\mid z_{t}$ $\sim\mathcal{N}\left(\rho\sigma_{\eta}z_{t},(1-\rho^{2})\sigma_{\eta}^{2}\right),$ (33)
where $\lambda_{t}\sim\mathcal{IG}(\nu/2,\nu/2)$, $\mu_{\lambda}=\nu/(\nu-2)$, and $\mathcal{TN}_{(0,\infty)}$ denotes a standard normal distribution truncated above zero. The constants
$c=\mathrm{E}[z_{0t}]=\sqrt{\frac{2}{\pi}},\quad\mathrm{Var}[z_{0t}]=1-c^{2}$
are used to ensure that $\epsilon_{t}$ is standardized, so that its conditional variance remains $\exp(h_{t})$.
The contemporaneous correlation between $\epsilon_{t}$ and $\eta_{t}$ is now a function of $\rho$, $\nu$, and $\delta$, and is given by
$\mathrm{Corr}[\epsilon_{t},\eta_{t}]=\sqrt{\frac{1-\delta^{2}}{(1-c^{2}\delta^{2})\mu_{\lambda}}}\cdot\mathrm{E}[\sqrt{\lambda_{t}}]\cdot\rho,$ (34)
where $\mathrm{E}[\sqrt{\lambda_{t}}]$ is defined in equation (24). Setting $\delta=0$ recovers the RSV-T model. Moreover, as $\nu\to\infty$ or $\lambda_{t}=1$, the model simplifies to the RSV-AZ-SN model with a skew-normal distribution.
The shape of the distribution of $\epsilon_{t}$ is governed jointly by $\delta$ and $\nu$, which respectively control skewness and tail thickness. From the general moment expression
$\mathrm{E}[\lambda_{t}^{m/2}]=\left(\frac{\nu}{2}\right)^{m/2}\cdot\frac{\Gamma\left(\frac{\nu-m}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)},\quad\nu>m,$ (35)
we derive the third and fourth moments of $\epsilon_{t}$ as:
$\mathrm{E}[\epsilon_{t}^{3}]$ $=\frac{4-\pi}{2}\cdot\frac{c^{3}\delta^{3}}{(1-c^{2}\delta^{2})^{3/2}}\cdot\left(\frac{\nu-2}{2}\right)^{3/2}\cdot\frac{\Gamma\left(\frac{\nu-3}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)},$ (36)
$\mathrm{E}[\epsilon_{t}^{4}]$ $=\left(3+2(\pi-3)\cdot\frac{c^{4}\delta^{4}}{(1-c^{2}\delta^{2})^{2}}\right)\cdot\frac{\nu-2}{\nu-4}.$ (37)
As with the GH skew-$t$ distribution, the skewness and kurtosis of the return distribution are jointly influenced by the parameters. While $\nu$ governs tail behavior, $\delta$ introduces asymmetry.
Figure 2 illustrates the impact of these parameters on the shape of $\epsilon_{t}$. Panel (i) shows that increasing negative values of $\delta$ lead to stronger left-skewness, with $\delta=0$ yielding a symmetric $t$-distribution. Panel (ii) demonstrates that larger values of $\nu$ reduce tail thickness and approximate the normal distribution. Notably, even when $\nu\to\infty$, the distribution remains skewed unless $\delta=0$, in which case the skew-normal distribution is recovered.
Figure 2: Simulated density of the standardized Azzalini skew-$t$ distribution in equation (30). (i) $\delta=0,-0.6,-0.9$ with $\nu=10$ fixed. (ii) $\nu=15,10,5$ with $\delta=-0.9$ fixed.
2.7 RSV Model with Fernández–Steel Skew-$t$ Distribution
Following the approach of Fernández and Steel (1995), we define the Fernández–Steel (FS) skew-$t$ density as
$p_{T}(w\mid\gamma,\nu)=\frac{2}{\gamma+\gamma^{-1}}\left\{f_{T}\left(\frac{w}{\gamma}\mid\nu\right)\mathbbm{1}\{w\geq 0\}+f_{T}\left(\gamma w\mid\nu\right)\mathbbm{1}\{w<0\}\right\},\quad\gamma>0,$ (38)
where $f_{T}(\cdot\mid\nu)$ denotes the probability density function of the standard Student’s $t$ distribution:
$f_{T}(\tilde{w}\mid\nu)=c_{\nu}\left(1+\frac{\tilde{w}^{2}}{\nu}\right)^{-\frac{\nu+1}{2}},\quad c_{\nu}=\frac{\Gamma\left(\frac{\nu+1}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\Gamma\left(\frac{1}{2}\right)\sqrt{\nu}},\quad\nu>2.$ (39)
Here, $\mathbbm{1}\{\cdot\}$ denotes the indicator function.
The mean and variance of $w\sim p_{T}(w\mid\gamma,\nu)$ are:
$\mu_{*}$ $=\mathrm{E}[w]=M_{1}(\gamma-\gamma^{-1}),$ (40)
$\sigma_{*}^{2}$ $=\mathrm{Var}[w]=M_{2}\cdot\frac{\gamma^{3}+\gamma^{-3}}{\gamma+\gamma^{-1}}-\left\{M_{1}(\gamma-\gamma^{-1})\right\}^{2},$ (41)
where
[…]
To obtain a standardized version of the FS skew-$t$ density, we define:
$w_{*}=\frac{w-\mu_{*}}{\sigma_{*}},$
so that $\mathrm{E}[w_{*}]=0$ and $\mathrm{Var}[w_{*}]=1$. The corresponding standardized density becomes:
$q_{T}(w_{*}\mid\gamma,\nu)$ $=\frac{2\sigma_{*}}{\gamma+\gamma^{-1}}\left\{f_{T}\left(\frac{\sigma_{*}w_{*}+\mu_{*}}{\gamma}\mid\nu\right)\mathbbm{1}\left\{w_{*}\geq-\frac{\mu_{*}}{\sigma_{*}}\right\}\right.$
$\quad\left.+f_{T}\left(\gamma(\sigma_{*}w_{*}+\mu_{*})\mid\nu\right)\mathbbm{1}\left\{w_{*}<-\frac{\mu_{*}}{\sigma_{*}}\right\}\right\}.$ (44)
Replacing $f_{T}$ with the standard normal density $f_{N}$ in the above expression yields the standardized FS skew-normal density $q_{N}(\cdot\mid\gamma)$.
We now define the RSV-FS-ST model by assuming:
$\epsilon_{t}$ $\sim q_{T}(\epsilon_{t}\mid\gamma,\nu),$ (45)
$\eta_{t}\mid\epsilon_{t}$ $\sim\mathcal{N}(\rho\sigma_{\eta}\epsilon_{t},(1-\rho^{2})\sigma_{\eta}^{2}),$ (46)
where $q_{T}(\cdot\mid\gamma,\nu)$ is the standardized FS skew-$t$ density. Replacing $q_{T}$ with $q_{N}$ yields the RSV-FS-SN model.
The skewness of the distribution is governed by the parameter $\gamma$, where $\gamma=1$ corresponds to a symmetric $t$-distribution. If $\gamma<1$, the distribution becomes left-skewed, while $\gamma>1$ results in right-skewness. The parameter $\nu$ controls tail heaviness.
Figure 3 illustrates how the shape of the distribution varies with $\gamma$ and $\nu$. Panel (i) shows that decreasing $\gamma$ induces stronger negative skewness, with $\gamma=1$ yielding symmetry. Panel (ii) demonstrates that increasing $\nu$ reduces tail thickness. Notably, as $\nu\to\infty$, the distribution converges to a skew-normal form when $\gamma\neq 1$, and to a standard normal when $\gamma=1$.
Figure 3: Density of the standardized FS skew-$t$ distribution in equation (44). (i) $\gamma=1,0.8,0.6$ with $\nu=10$. (ii) $\nu=20,15,10$ with $\gamma=0.5$.
3 Bayesian Estimation Methodology
Each RSV model introduced in Section 2 can be formulated as a nonlinear Gaussian state space model with a large number of latent variables. This structure renders the likelihood function intractable in closed form, and consequently makes maximum likelihood estimation computationally difficult. To overcome this, we adopt a Bayesian estimation framework via MCMC simulation, following established approaches in the literature.
We assign prior distributions to the common model parameters $(\mu,\phi,\sigma_{\eta}^{2},\rho,\xi,\sigma_{u}^{2})$ as follows:
$\mu$ $\sim\mathcal{N}(m_{\mu},s_{\mu}^{2}),\quad\frac{\phi+1}{2}\sim\mathcal{B}(a_{\phi 0},b_{\phi 0}),$ (47)
$\frac{\rho+1}{2}$ $\sim\mathcal{B}(a_{\rho 0},b_{\rho 0}),\quad\sigma_{\eta}^{2}\sim\mathcal{IG}\left(\frac{n_{\eta}}{2},\frac{S_{\eta}}{2}\right),$ (48)
$\xi$ $\sim\mathcal{N}(m_{\xi},s_{\xi}^{2}),\quad\sigma_{u}^{2}\sim\mathcal{IG}\left(\frac{n_{u}}{2},\frac{S_{u}}{2}\right),$ (49)
where $\mathcal{B}(\cdot,\cdot)$ and $\mathcal{IG}(\cdot,\cdot)$ denote the beta and inverse-gamma distributions, respectively. The hyperparameters $(m_{\mu},s_{\mu},a_{\phi 0},b_{\phi 0},a_{\rho 0},b_{\rho 0},n_{\eta},S_{\eta},m_{\xi},s_{\xi},n_{u},S_{u})$ are chosen to reflect prior beliefs or to be weakly informative when such prior knowledge is unavailable.
For notational convenience, let $\bm{y}=(y_{1},\ldots,y_{n})^{\prime}$, $\bm{x}=(x_{1},\ldots,x_{n})^{\prime}$, and $\bm{h}=(h_{1},\ldots,h_{n})^{\prime}$. Denote the parameter vector by $\bm{\theta}$, and let $\pi(\bm{\theta})$ be its joint prior density. The joint likelihood of $(\bm{y},\bm{x},\bm{h})$ given $\bm{\theta}$ is denoted by $f(\bm{y},\bm{x},\bm{h}\mid\bm{\theta})$.
In the following, we detail the MCMC sampling algorithms specifically for the RSV-AZ-ST and RSV-FS-ST models. [^2] See Takahashi et al. (2009) and Takahashi et al. (2016) for the algorithms used in the RSV-N and RSV-GH-ST models, respectively. We also outline the procedure for computing one-step-ahead forecasts of volatility and returns, from which we derive one-step-ahead VaR and ES.
3.1 RSV-AZ-ST Model
Let $\bm{\theta}=(\mu,\phi,\rho,\sigma_{\eta}^{2},\delta,\xi,\sigma_{u}^{2},\nu)^{\prime}$ denote the full vector of model parameters, and define the latent variable vectors $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{n})^{\prime}$ and $\bm{z}_{0}=(z_{01},\ldots,z_{0n})^{\prime}$. For the prior distributions of the model-specific parameters $\delta$ and $\nu$, we assume:
$\frac{\delta+1}{2}$ $\sim\mathcal{B}(a_{\delta 0},b_{\delta 0}),\quad\nu\sim\mathcal{G}(n_{\nu 0},S_{\nu 0}),$ (50)
where $\mathcal{G}(\cdot,\cdot)$ denotes the gamma distribution.
The MCMC algorithm proceeds as follows:
  1. 1.
    Initialize $\bm{h}$, $\bm{\theta}$, $\bm{z}_{0}$, and $\bm{\lambda}$.
  2. 2.
    Sample $\mu\mid\bm{\theta}_{-\mu},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  3. 3.
    Sample $\phi\mid\bm{\theta}_{-\phi},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  4. 4.
    Jointly sample $(\rho,\sigma_{\eta}^{2})\mid\bm{\theta}_{-(\rho,\sigma_{\eta}^{2})},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  5. 5.
    Sample $\delta\mid\bm{\theta}_{-\delta},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  6. 6.
    Sample $\xi\mid\bm{\theta}_{-\xi},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  7. 7.
    Sample $\sigma_{u}^{2}\mid\bm{\theta}_{-\sigma_{u}^{2}},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  8. 8.
    Sample $\nu\mid\bm{\theta}_{-\nu},\bm{h},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  9. 9.
    Sample $\bm{z}_{0}\mid\bm{\theta},\bm{h},\bm{\lambda},\bm{x},\bm{y}$.
  10. 10.
    Sample $\bm{h}\mid\bm{\theta},\bm{z}_{0},\bm{\lambda},\bm{x},\bm{y}$.
  11. 11.
    Sample $\bm{\lambda}\mid\bm{\theta},\bm{h},\bm{z}_{0},\bm{x},\bm{y}$.
  12. 12.
    Repeat steps 2–11 for a sufficiently large number of iterations.
Details of each sampling step, including full conditional distributions and implementation strategies (e.g., Gibbs sampling or Metropolis–Hastings updates), are provided in Appendix A.1.
3.2 RSV-FS-ST Model
Let $\bm{\theta}=(\mu,\phi,\rho,\sigma_{\eta}^{2},\gamma,\xi,\sigma_{u}^{2},\nu)^{\prime}$ denote the full set of model parameters. For the model-specific parameters $\gamma$ and $\nu$, we assign the following prior distributions:
$\gamma\sim\mathcal{G}(n_{\gamma 0},S_{\gamma 0}),\quad\nu\sim\mathcal{G}(n_{\nu 0},S_{\nu 0}).$ (51)
The MCMC algorithm proceeds as follows:
  1. 1.
    Initialize $\bm{h}$ and $\bm{\theta}$.
  2. 2.
    Sample $\mu\mid\bm{\theta}_{-\mu},\bm{h},\bm{x},\bm{y}$.
  3. 3.
    Sample $\phi\mid\bm{\theta}_{-\phi},\bm{h},\bm{x},\bm{y}$.
  4. 4.
    Sample $\rho\mid\bm{\theta}_{-\rho},\bm{h},\bm{x},\bm{y}$.
  5. 5.
    Sample $\sigma_{\eta}^{2}\mid\bm{\theta}_{-\sigma_{\eta}^{2}},\bm{h},\bm{x},\bm{y}$.
  6. 6.
    Sample $\xi\mid\bm{\theta}_{-\xi},\bm{h},\bm{x},\bm{y}$.
  7. 7.
    Sample $\sigma_{u}^{2}\mid\bm{\theta}_{-\sigma_{u}^{2}},\bm{h},\bm{x},\bm{y}$.
  8. 8.
    Sample $\gamma\mid\bm{\theta}_{-\gamma},\bm{h},\bm{x},\bm{y}$.
  9. 9.
    Sample $\nu\mid\bm{\theta}_{-\nu},\bm{h},\bm{x},\bm{y}$.
  10. 10.
    Sample $\bm{h}\mid\bm{\theta},\bm{x},\bm{y}$.
  11. 11.
    Repeat steps 2–10 for a sufficiently large number of iterations.
The full conditional distributions and implementation details for each step are provided in Appendix A.2.
3.3 One-day-ahead Forecast
To obtain one-day-ahead forecasts of financial returns and volatilities, we utilize the predictive distribution within each state-space model. Let $\bm{\theta}^{(i)}$ and $\bm{h}^{(i)}$ represent the $i$th sample of parameters and latent log-volatilities in the MCMC simulation, respectively. The one-step-ahead predictive samples are then generated as follows:
  1. 1.
    Generate $h_{n+1}^{(i)}\mid\bm{x},\bm{y},\bm{h}^{(i)},\bm{\theta}^{(i)}\sim N(\mu_{n+1}^{(i)},(\sigma_{n+1}^{(i)})^{2})$ where
    $\mu_{n+1}^{(i)}$ $=\mu^{(i)}+\phi^{(i)}h_{n}^{(i)}-\mu^{(i)}+\rho^{(i)}\sigma_{\eta}^{(i)}y_{n}\exp\left(-\frac{h_{n}^{(i)}}{2}\right),$ (52)
    $(\sigma_{n+1}^{(i)})^{2}$ $=(1-(\rho^{(i)})^{2})(\sigma_{\eta}^{(i)})^{2}.$ (53)
  2. 2.
    Generate $\epsilon_{n+1}^{(i)}$ from the distribution corresponding to the chosen RSV model given $\bm{\theta}^{(i)}$ and $\bm{h}^{(i)}$.
  3. 3.
    Compute $y_{n+1}^{(i)}$ as
    $y_{n+1}^{(i)}=\epsilon_{n+1}^{(i)}\exp\left(-\frac{h_{n+1}^{(i)}}{2}\right).$ (54)
Repetition of the above procedure for $M$ times allows us to obtain the generated samples $\{h_{n+1}^{(i)}\}_{i=1}^{M}$ and $\{y_{n+1}^{(i)}\}_{i=1}^{M}$.
One-day-ahead quantile forecasts, such as VaR and ES, can be derived from the distribution of $\{y_{n+1}^{(i)}\}_{i=1}^{M}$ for each model. The one-day-ahead VaR forecast at time $n+1$ at level $\alpha$, denoted by $\mbox{VaR}_{n+1}(\alpha)$, is defined as
$\mbox{Pr}(y_{n+1}<\mbox{VaR}_{n+1}(\alpha)\mid\mathcal{I}_{n})=\alpha,$ (55)
where $\mathcal{I}_{n}$ denotes the information set available at time $n$. The corresponding ES forecast, denoted by $\mbox{ES}_{n+1}(\alpha)$, is given as
$\mbox{ES}_{n+1}(\alpha)=E[y_{n+1}\mid y_{n+1}<\mbox{VaR}_{n+1}(\alpha),\mathcal{I}_{n}].$ (56)
The one-day-ahead VaR and ES can be obtained as the $(1-\alpha)$th quantile and conditional average of $\{y_{n+1}^{(i)}\}_{i=1}^{M}$, respectively.
We evaluate the obtained forecasts for volatility, VaR, and ES using appropriate loss or scoring functions. To evaluate the volatility forecasts, we compute the Gaussian quasi-likelihood (QLIKE) loss function, defined as:
$L_{QLIKE}(x,f)=\frac{x}{f}-\log\frac{x}{f}-1,$
where $x$ and $f$ represent a volatility proxy and a volatility forecast, respectively. The QLIKE loss function is robust, as defined by Patton (2011), meaning that it provides consistent rankings regardless of whether the ranking is based on true volatility or a conditionally unbiased volatility proxy. Furthermore, Patton and Sheppard (2009) demonstrated its superior power compared to the mean squared error (MSE), another robust loss function, in the predictive precision test proposed by Diebold and Mariano (2002).
For evaluating the forecasts of VaR and ES, we employ the joint loss function introduced by Fissler and Ziegel (2016) (FZ loss). Following the approach of Patton et al. (2019), we utilize a specific form of the FZ loss function, referred to as the FZ0 loss function, which is expressed as:
$L_{FZ0}(y,v,e;\alpha)=-\frac{1}{\alpha e}\mathbbm{1}\{y\leq v\}(v-y)+\frac{v}{e}+\log(-e)-1,$ (57)
where $y$, $v$, and $e$ represent a return, VaR, and ES, respectively. Patton et al. (2019) demonstrated that the FZ0 loss function is unique in producing loss differences that are homogeneous of degree zero.
While average losses derived from the aforementioned loss functions offer initial insights into the forecast performance of the models in contention, they do not indicate if the differences in losses are statistically significant. To ascertain this, we utilize the conditional and unconditional predictive ability tests by Giacomini and White (2006), hereinafter referred to as GW tests, on the loss differences. The GW tests are particularly pertinent to this paper’s objectives. This is because they accommodate the rolling window methods adopted in Section 4, enabling a unified assessment of both nested and non-nested models, including the RSV models described in Section 2.
For the unconditional predictive ability, the GW test statistic conforms to that proposed by Diebold and Mariano (2002) and is, under the null hypothesis, asymptotically standard normally distributed. For the conditional predictive ability, the GW test defines the null hypothesis as
$H_{0,h}:\mathrm{E}[\Delta L_{t+1}\mid\mathcal{I}_{t}]=0,\quad t=n,n+1,\ldots,n+n_{f}-1,$ (58)
where $\Delta L_{t}$ is the loss difference between two models at forecast date $t$, and $\mathcal{I}_{t}$ is the information set at time $t$.
Given a test function $\mathbbm{h}_{t}$, a $q\times 1$ vector measurable with respect to $\mathcal{I}_{t}$ (Stinchcombe and White, 1998), the null hypothesis translates to
$H_{0,h}:\mathrm{E}[\mathbbm{h}_{t}\Delta L_{t+1}]=0,\quad t=n,n+1,\ldots,n+n_{f}-1.$ (59)
Defining $Z_{t+1}=\mathbbm{h}_{t}\Delta L_{t+1}$ and $\bar{Z}_{n_{f}}=n_{f}^{-1}\sum_{t=n}^{n+n_{f}-1}Z_{t+1}$, the alternative hypothesis becomes
$H_{A,h}:\mathrm{E}[\bar{Z}_{n_{f}}^{\prime}]\mathrm{E}[\bar{Z}_{n_{f}}]\geq d>0,\quad\mbox{for sufficiently large $n_{f}$}.$ (60)
We employ the Wald-type test statistic
$T_{n_{f}}=n\bar{Z}_{n_{f}}^{\prime}\hat{\Omega}_{n_{f}}^{-1}\bar{Z}_{n_{f}},$ (61)
where $\hat{\Omega}_{n_{f}}=n_{f}^{-1}\sum_{t=n}^{n+n_{f}-1}Z_{t+1}Z_{t+1}^{\prime}$ consistently estimates the variance of $Z_{t+1}$. Under the null, $T_{n_{f}}$ follows a chi-square distribution with $q$ degrees of freedom, $\chi_{q}^{2}$. At the significance level $p$, the null hypothesis is rejected if $T_{n_{f}}>\chi_{q,1-p}^{2}$, where $\chi_{q,1-p}^{2}$ is the $(1-p)$ quantile of $\chi_{q}^{2}$.
In practice, the selection of $\mathbbm{h}_{t}$ should discern between the forecast performances of the models. Focusing on one-day-ahead forecasts, we follow the specifications of Takahashi et al. (2023) and define $\mathbbm{h}_{t}=(1,\Delta L_{t})^{\prime}$ for $t=n,n+1,\ldots,n+n_{f}-1$. If the null is rejected, it implies that the lagged loss differences, $\Delta L_{t}$, can help predict the subsequent loss differences, $\Delta L_{t+1}$. Using $\hat{b}$ to represent the regression coefficient of $\Delta L_{t+1}$ on $\mathbbm{h}_{t}$, the predicted loss differences $\{\hat{b}^{\prime}\mathbbm{h}_{t}\}_{t=n}^{n+n_{f}-1}$ serve as indicators for assessing model performances across different time points. In Section 4, we measure relative performance by the frequency with which one model forecasts larger losses than the other, as expressed by
$I_{n_{f},0}=\frac{1}{n_{f}}\sum_{t=n}^{n+n_{f}-1}I(\hat{b}^{\prime}\mathbbm{h}_{t}>0).$ (62)
4 Empirical Study
4.1 Data and Descriptive Statistics
We estimate the SV and RSV models described in Section 2 using daily (close-to-close) returns and RVs for two major stock indices: the DJIA and the N225. Following Liu et al. (2015), we adopt the 5-minute RV estimator, computed during trading hours only, among several available alternatives. The DJIA data are obtained from the Oxford-Man Institute’s Realized Library, [^3] The website https://realized.oxford-man.ox.ac.uk/ is no longer accessible. while the N225 data are constructed from the Nikkei NEEDS-TICK dataset. [^4] See Ubukata and Watanabe (2014) for details on the construction of the N225 dataset.
The sample period spans from June 2009 to September 2019 for both indices. Specifically, the DJIA sample covers 2,596 trading days from June 1, 2009, to September 27, 2019, and the N225 sample includes 2,532 trading days from June 1, 2009, to September 30, 2019.
Figure 4 presents time series plots and histograms of daily returns for the DJIA and N225. Both return series exhibit considerable variation around zero and show frequent large negative returns, indicating negatively skewed distributions.
DJIA N225
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Time series plots (top) and histograms (bottom) of daily returns (in percentage points) for the DJIA and N225.
Table 1 reports the descriptive statistics for the daily returns. The mean return is significantly different from zero for the DJIA, but not for the N225. However, since the means are negligible in magnitude, we do not demean the series in the subsequent analysis. The $p$-values from the Ljung–Box statistic (Ljung and Box, 1978), adjusted for heteroskedasticity as in Diebold (1988), do not reject the null of no autocorrelation up to 10 lags in either series. This allows us to estimate the models directly using raw returns.
Both return series are characterized by significantly negative skewness and high kurtosis, suggesting leptokurtic distributions—a stylized fact in financial returns. The Jarque–Bera (JB) statistic confirms that normality is strongly rejected for both series. These empirical features motivate our use of skewed-$t$ distributions for the return innovations, as discussed in Section 2.
Table 1: Descriptive statistics of daily returns for the DJIA and N225
Mean SD Skew Kurt Min Max JB LB
DJIA $0.044$ $0.895$ $-0.448$ $6.675$ $-5.562$ $4.857$ $0.00$ $0.52$
$(0.018)$ $(0.048)$ $(0.096)$
N225 $0.033$ $1.321$ $-0.544$ $8.101$ $-11.153$ $7.426$ $0.00$ $0.55$
$(0.026)$ $(0.049)$ $(0.097)$
  • Notes: Standard errors are shown in parentheses. JB refers to the $p$-value of the Jarque–Bera test. LB denotes the $p$-value of the Ljung and Box (1978) statistic, adjusted as in Diebold (1988).
Figure 5 shows the time series and histograms of the 5-minjte RVs and their logarithmic transformations for both indices. Both series exhibit strong temporal clustering, high persistence, and occasional sharp spikes. [^5] For the DJIA, the largest RV spike occurred on August 24, 2015, amid heightened market turbulence (see The New York Times: https://www.nytimes.com/2015/08/25/business/dealbook/daily-stock-market-activity.html). For the N225, the most notable spike on March 15, 2011, reflects the aftermath of the Great East Japan Earthquake on March 11, 2011. These spikes lead to positively skewed distributions in the log-RVs.
DJIA N225
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Time series plots of 5-minute RVs (top), their logarithms (middle), and corresponding histograms (bottom) for the DJIA and N225.
Table 2 summarizes the descriptive statistics of the log-RVs. The Ljung–Box test strongly rejects the null of no autocorrelation, affirming the presence of volatility clustering in both markets. The distributions are positively skewed and leptokurtic. Consequently, the JB test rejects the normality of the log-RVs. This finding is inconsistent with the normality assumption for the error term $u_{t}$ in equation (16). While we maintain this assumption for tractability, allowing for non-normality in future model extensions would be a valuable research direction.
Table 2: Descriptive statistics of the logarithms of 5-minute RVs for the DJIA and N225
Mean SD Skew Kurt Min Max JB LB
DJIA $-0.987$ $1.048$ $0.317$ $3.177$ $-3.953$ $4.087$ $0.00$ $0.00$
$(0.021)$ $(0.048)$ $(0.096)$
N225 $-0.858$ $0.844$ $0.589$ $4.219$ $-3.321$ $3.258$ $0.00$ $0.00$
$(0.017)$ $(0.049)$ $(0.097)$
  • Notes: See Table 1 for additional details.
4.2 Competing Models
We evaluate the forecasting performance of the following RSV models, which differ in the distributional assumption for $\epsilon_{t}$:
  1. 1.
    RSV-N: $\epsilon_{t}$ follows a normal distribution.
  2. 2.
    RSV-T: $\epsilon_{t}$ follows a Student’s $t$ distribution.
  3. 3.
    RSV-GH-ST: $\epsilon_{t}$ follows the GH skew-$t$ distribution.
  4. 4.
    RSV-AZ-SN: $\epsilon_{t}$ follows the Azzalini skew normal distribution.
  5. 5.
    RSV-AZ-ST: $\epsilon_{t}$ follows the Azzalini skew-$t$ distribution.
  6. 6.
    RSV-FS-SN: $\epsilon_{t}$ follows the Fernández–Steel skew normal distribution.
  7. 7.
    RSV-FS-ST: $\epsilon_{t}$ follows the Fernández–Steel skew-$t$ distribution.
The prior distributions for the common parameters are specified as:
$\mu$ $\sim\mathcal{N}(0,100),\quad\frac{\phi+1}{2}\sim\mathcal{B}(1,1),\quad\frac{\rho+1}{2}\sim\mathcal{B}(1,1),\quad\sigma_{\eta}^{2}\sim\mathcal{IG}(0.05,0.05),$ (63)
$\xi$ $\sim\mathcal{N}(0,10),\quad\sigma_{u}^{2}\sim\mathcal{IG}(2.5,0.1).$ (64)
For the additional distributional parameters, we set:
$\nu\sim\mathcal{G}(5,0.5),\quad\beta\sim\mathcal{N}(0,1),\quad\frac{\delta+1}{2}\sim\mathcal{B}(1,1),\quad\gamma\sim G(1,1).$ (65)
In addition, to isolate the contribution of RV, we estimate standard SV models with the same innovation distributions as in the RSV specifications. These include SV-N, SV-T, SV-GH-ST, SV-AZ-SN, SV-AZ-ST, SV-FS-SN, and SV-FS-ST, mirroring the distributional assumptions used in the RSV models. This setup allows us to directly assess the added value of incorporating RV in terms of predictive performance.
To further benchmark the RSV models, we also consider GARCH-type models, which are widely used to capture time-varying volatility, including ARCH (Engle, 1982), GARCH (Bollerslev, 1986), and their variants. Among these, we employ the EGARCH model introduced by Nelson (1991). Furthermore, we include the REGARCH model, a generalized version of the RGARCH framework that incorporates RV measures. See Hansen et al. (2012); Hansen and Huang (2016) for foundational work, and Takahashi et al. (2023) for an overview.
Both EGARCH and REGARCH models are estimated via the quasi-maximum likelihood (QML) method under standard normal and Student’s $t$ innovations. One-day-ahead forecasts of volatility, VaR, and ES are computed using the estimated parameters. For Bayesian approaches to RGARCH-type models and their applications to volatility and tail risk forecasting, see Chen et al. (2023).
4.3 Out-of-sample Forecasting
Following the approach of Takahashi et al. (2016), we implement a rolling window estimation procedure to evaluate the out-of-sample performance of each model. The window size is kept fixed throughout the forecasting period. For the DJIA, we use a window of 1,993 observations, generating forecasts from May 1, 2017, to September 27, 2019. For the N225, the window consists of 1,942 observations, with forecast dates ranging from May 1, 2017, to September 30, 2019. After each estimation step, we produce one-day-ahead forecasts of volatility, VaR, and ES.
At each forecast point, we draw 15,000 predictive samples from the posterior predictive distribution. For the SV and RSV models, we compute both the posterior means and medians of the volatility forecasts. [^6] We report the median rather than the mean for volatility forecasts of SV models, as their posterior distributions occasionally exhibit heavy tails, which can disproportionately inflate the means. Additionally, we calculate predictive quantiles of the return distribution to obtain VaR and ES estimates. This procedure yields 603 forecasts for the DJIA and 590 for the N225, covering the period from early May 2017 to late September 2019.
4.3.1 Volatility Forecasts
To evaluate the accuracy of volatility forecasts, we compute the average loss using the QLIKE loss function, as described in Section 3.3. The QLIKE loss is known for its robustness, yielding model rankings that align closely with those based on latent volatility, provided the proxy is conditionally unbiased. Simulation evidence from Patton and Sheppard (2009) and empirical findings by Hansen and Lunde (2005) and Patton (2011) suggest that QLIKE has greater statistical power than MSE to discriminate between models.
To mitigate market microstructure noise, we employ multiple volatility proxies: realized kernel (RK) with a flat-top Tukey-Hanning2 kernel (Barndorff-Nielsen et al., 2008), bipower variation (BV) (Barndorff-Nielsen and Shephard, 2004), and median realized volatility (Med) (Andersen et al., 2012), in addition to the standard 5-minute realized volatility (RV5).
To correct for biases due to non-trading hours, we adopt the adjustment method proposed by Hansen and Lunde (2005), as given in equation (9), using the following correction factor:
$c_{HL}=\frac{\sum_{s=t-n+1}^{t}(y_{s}-\bar{y})^{2}}{\sum_{s=t-n+1}^{t}x_{s}},\quad\bar{y}=\frac{1}{n}\sum_{s=t-n+1}^{t}y_{s},$ (66)
where $n$ is the window size and $x_{s}$ denotes the volatility proxy at time $s$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Log-scale volatility forecasts together with the adjusted RV5 for DJIA: Overall period (top) and sub-periods (bottom).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Log-scale volatility forecasts together with the adjusted RV5 for N225: Overall period (top) and sub-periods (bottom).
Figures 6 and 7 display the log-scale volatility forecasts alongside the adjusted RV5 series for the DJIA and N225. For the DJIA, SV model forecasts—especially those from the SV-GH-ST specification—exhibit pronounced volatility spikes. Across both indices, forecast patterns are more homogeneous within model families (e.g., SV, RSV, EGARCH, REGARCH) than across different distributional assumptions, indicating that model structure plays a more dominant role than distributional form.
Table 3: QLIKE scores of volatility forecasts for DJIA and N225
DJIA N225 Average
RV5 RK BV Med RV5 RK BV Med Score Rank
SV-N $0.349$ $0.532$ $0.587$ $0.641$ $0.176$ $0.177$ $0.169$ $0.169$ $0.350$ $14.8$
SV-T $0.329$ $0.502$ $0.563$ $0.620$ $0.189$ $0.192$ $0.178$ $0.172$ $0.343$ $15.4$
SV-AZ-SN $0.362$ $0.557$ $0.617$ $0.680$ $0.183$ $0.183$ $0.174$ $0.174$ $0.366$ $16.4$
SV-AZ-ST $0.337$ $0.519$ $0.580$ $0.646$ $0.189$ $0.191$ $0.179$ $0.174$ $0.352$ $16.0$
SV-FS-SN $0.273^{*}$ $0.392^{*}$ $0.446^{*}$ $0.502^{*}$ $0.165$ $0.170$ $0.159$ $0.146$ $0.282$ $7.2$
SV-FS-ST $0.267^{*}$ $0.378^{*}$ $0.432^{*}$ $0.487^{*}$ $0.172$ $0.178$ $0.165$ $0.148$ $0.278$ $7.8$
SV-GH-ST $0.346$ $0.522$ $0.582$ $0.657$ $0.188$ $0.189$ $0.181$ $0.174$ $0.355$ $16.5$
RSV-N $0.245^{*}$ $0.452^{*}$ $0.486^{*}$ $0.567^{*}$ $0.043^{*}$ $0.047^{*}$ $0.047^{*}$ $0.043^{*}$ $0.241$ $6.0$
RSV-T $0.231^{*}$ $0.435^{*}$ $0.471^{*}$ $0.551^{*}$ $0.043^{*}$ $0.048$ $0.047^{*}$ $0.041^{*}$ $0.234$ $4.6$
RSV-AZ-SN $0.240^{*}$ $0.449^{*}$ $0.485^{*}$ $0.566^{*}$ $0.043^{*}$ $0.048^{*}$ $0.048^{*}$ $0.043^{*}$ $0.240$ $6.1$
RSV-AZ-ST $0.227^{*}$ $0.425^{*}$ $0.462^{*}$ $0.540^{*}$ $0.043^{*}$ $0.049$ $0.048^{*}$ $0.041^{*}$ $0.229$ $4.1$
RSV-FS-SN $0.252^{*}$ $0.471$ $0.505^{*}$ $0.589$ $0.042^{*}$ $0.046^{*}$ $0.047^{*}$ $0.045^{*}$ $0.250$ $7.2$
RSV-FS-ST $0.242^{*}$ $0.452^{*}$ $0.487^{*}$ $0.569^{*}$ $0.042^{*}$ $0.047^{*}$ $0.047^{*}$ $0.043^{*}$ $0.241$ $6.0$
RSV-GH-ST $0.223^{*}$ $0.467^{*}$ $0.498^{*}$ $0.577^{*}$ $0.044$ $0.047^{*}$ $0.046^{*}$ $0.040^{*}$ $0.243$ $5.9$
EGARCH-N $0.327$ $0.441^{*}$ $0.510^{*}$ $0.575^{*}$ $0.167$ $0.175$ $0.162$ $0.141$ $0.312$ $11.0$
EGARCH-T $0.322$ $0.437^{*}$ $0.505^{*}$ $0.558^{*}$ $0.183$ $0.191$ $0.176$ $0.155$ $0.316$ $12.0$
REGARCH-N $0.248^{*}$ $0.428^{*}$ $0.476^{*}$ $0.556^{*}$ $0.058$ $0.063$ $0.060$ $0.046^{*}$ $0.242$ $7.6$
REGARCH-T $0.245^{*}$ $0.424^{*}$ $0.472^{*}$ $0.551^{*}$ $0.056$ $0.062$ $0.059$ $0.045^{*}$ $0.239$ $6.4$
  • Notes: An asterisk () indicates that the model belongs to the 90% MCS. The average columns report the mean QLIKE scores and ranks across the two indices.
Table 3 reports QLIKE scores across all models and volatility proxies. An asterisk () denotes inclusion in the 90% model confidence set (MCS) of Hansen et al. (2011). [^7] The rolling window scheme satisfies the stationarity assumption required by the MCS bootstrap. We use 1,000 bootstrap replications with a block size of 10. See Hansen et al. (2011), Section 4.3. Overall, RSV and REGARCH models attain lower QLIKE scores than SV and EGARCH counterparts. For the DJIA, SV-FS-ST achieves the lowest scores under RK, BV, and Med. For the N225, RSV models dominate REGARCH, whereas the two perform comparably for the DJIA. On average, the RSV-AZ-ST model delivers the best performance across indices.
RV5 RK
Refer to caption Refer to caption
BV Med
Refer to caption Refer to caption
Figure 8: Cumulative QLIKE loss differences in comparison to the SV-N model for DJIA.
RV5 RK
Refer to caption Refer to caption
BV Med
Refer to caption Refer to caption
Figure 9: Cumulative QLIKE loss differences in comparison to the SV-N model for N225.
To assess time-varying performance, Figures 8 and 9 depict cumulative loss differences (CLDs) relative to SV-N. Positive values indicate superior performance. When RV5 is used as the volatility proxy, the RSV and REGARCH models consistently outperform the SV and EGARCH models across the entire forecast horizon for both the DJIA and N225, underscoring the value of incorporating RV in prediction. For the DJIA, however, when alternative proxies such as RK, BV, or Med are used, the SV-FS-SN and SV-FS-ST models exhibit improved performance after mid-2018, occasionally surpassing RSV and REGARCH. In contrast, for the N225, RSV and REGARCH models maintain superior performance across all proxies, with clearer distinctions emerging between the two model classes in late 2017.
Table 4: Summary of GW tests on QLIKE of volatility forecasts for DJIA and N225
DJIA
SV-N SV-T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N EG-T REG-N REG-T
SV-N $90$ $12$ $77$ $98$ $96$ $58$ $94$ $96$ $94$ $96$ $92$ $94$ $97$ $67$ $71$ $92$ $93$
SV-T $-2.3$ $4$ $9$ $92$ $98$ $14$ $93$ $96$ $93$ $96$ $90$ $94$ $96$ $50$ $56$ $93$ $94$
SV-AZ-SN $2.2$ $4.3$ $94$ $96$ $97$ $79$ $97$ $98$ $97$ $98$ $95$ $97$ $98$ $71$ $74$ $95$ $96$
SV-AZ-ST $-1.4$ $1.7$ $-4.7$ $92$ $94$ $26$ $94$ $96$ $94$ $96$ $92$ $94$ $97$ $56$ $63$ $94$ $94$
SV-FS-SN $-7.3$ $-5.4$ $-7.4$ $-5.8$ $84$ $2$ $65$ $75$ $69$ $78$ $57$ $66$ $80$ $10$ $13$ $68$ $69$
SV-FS-ST $-6.2$ $-6.1$ $-6.7$ $-6.0$ $-1.1$ $2$ $60$ $72$ $65$ $77$ $54$ $61$ $80$ $7$ $9$ $64$ $67$
SV-GH-ST $-0.2$ $1.1$ $-1.1$ $0.6$ $4.5$ $4.7$ $98$ $98$ $98$ $98$ $97$ $98$ $99$ $66$ $67$ $95$ $95$
RSV-N $-5.3$ $-4.6$ $-6.0$ $-5.1$ $-1.7$ $-1.3$ $-4.7$ $100$ $97$ $99$ 00 $99$ $99$ $15$ $20$ $59$ $63$
RSV-T $-6.0$ $-5.4$ $-6.7$ $-6.0$ $-2.5$ $-2.1$ $-5.4$ $-5.6$ $3$ $81$ $3$ $1$ $91$ $9$ $14$ $39$ $44$
RSV-AZ-SN $-5.4$ $-4.8$ $-6.1$ $-5.3$ $-1.9$ $-1.5$ $-4.9$ $-1.7$ $6.6$ $98$ $9$ $25$ $97$ $13$ $18$ $52$ $57$
RSV-AZ-ST $-6.0$ $-5.6$ $-6.7$ $-6.1$ $-2.7$ $-2.4$ $-5.4$ $-5.0$ $-2.5$ $-9.6$ $6$ $8$ $91$ $9$ $13$ $29$ $35$
RSV-FS-SN $-4.9$ $-4.1$ $-5.7$ $-4.7$ $-1.2$ $-0.8$ $-4.3$ $2.7$ $8.2$ $4.3$ $6.4$ $97$ $95$ $20$ $26$ $66$ $71$
RSV-FS-ST $-5.4$ $-4.7$ $-6.2$ $-5.3$ $-1.8$ $-1.4$ $-4.8$ $-1.1$ $8.5$ $1.5$ $5.8$ $-6.6$ $94$ $15$ $18$ $56$ $59$
RSV-GH-ST $-6.2$ $-5.8$ $-7.0$ $-6.4$ $-3.0$ $-2.6$ $-5.7$ $-6.2$ $-5.0$ $-9.9$ $-4.7$ $-7.4$ $-7.2$ $8$ $11$ $24$ $28$
EG-N $-1.2$ $-0.1$ $-1.8$ $-0.6$ $4.3$ $5.4$ $-0.9$ $3.8$ $4.5$ $4.0$ $4.7$ $3.4$ $3.9$ $4.9$ $68$ $89$ $92$
EG-T $-1.3$ $-0.4$ $-1.9$ $-0.8$ $3.1$ $4.2$ $-1.0$ $3.1$ $3.8$ $3.4$ $4.0$ $2.8$ $3.2$ $4.2$ $-0.7$ $86$ $89$
REG-N $-4.6$ $-4.3$ $-5.1$ $-4.4$ $-1.4$ $-1.1$ $-4.2$ $0.2$ $1.3$ $0.6$ $1.8$ $-0.3$ $0.4$ $2.1$ $-4.1$ $-3.6$ $92$
REG-T $-4.6$ $-4.3$ $-5.1$ $-4.4$ $-1.5$ $-1.2$ $-4.2$ $0.0$ $1.1$ $0.4$ $1.5$ $-0.5$ $0.2$ $1.7$ $-4.2$ $-3.7$ $-1.9$
N225
SV-N SV-T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N EG-T REG-N REG-T
SV-N $4$ 00 $3$ $56$ $50$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $58$ $28$ $100$ $100$
SV-T $1.3$ $83$ $66$ $99$ $88$ $81$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $89$ $56$ $100$ $100$
SV-AZ-SN $2.2$ $-0.8$ $15$ $72$ $56$ $17$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $70$ $40$ $100$ $100$
SV-AZ-ST $1.5$ $-0.1$ $1.0$ $96$ $77$ $69$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $85$ $52$ $100$ $100$
SV-FS-SN $-1.4$ $-1.7$ $-1.8$ $-1.8$ $4$ $4$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $33$ $9$ $100$ $100$
SV-FS-ST $-0.6$ $-1.5$ $-1.2$ $-1.6$ $1.4$ $1$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $74$ $13$ $100$ $100$
SV-GH-ST $1.9$ $-0.1$ $0.8$ $-0.1$ $3.0$ $2.1$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $76$ $62$ $100$ $100$
RSV-N $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.2$ $-6.7$ $1$ 00 $1$ $100$ $92$ $3$ 00 00 00 00
RSV-T $-5.8$ $-4.8$ $-5.6$ $-5.0$ $-6.8$ $-6.1$ $-6.6$ $2.5$ $100$ $2$ $100$ $100$ 00 00 00 00 00
RSV-AZ-SN $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.1$ $-6.7$ $2.8$ $-1.6$ 00 $99$ $99$ $1$ 00 00 00 00
RSV-AZ-ST $-5.8$ $-4.8$ $-5.6$ $-5.0$ $-6.8$ $-6.1$ $-6.6$ $2.3$ $0.5$ $1.5$ $100$ $100$ 00 00 00 00 00
RSV-FS-SN $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.1$ $-6.7$ $-1.9$ $-2.4$ $-2.6$ $-2.3$ 00 $2$ 00 00 00 00
RSV-FS-ST $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.1$ $-6.7$ $-0.8$ $-3.0$ $-1.9$ $-2.8$ $1.3$ $1$ 00 00 00 00
RSV-GH-ST $-5.8$ $-4.7$ $-5.6$ $-5.0$ $-6.8$ $-6.1$ $-6.6$ $4.6$ $5.8$ $4.1$ $4.9$ $3.5$ $4.7$ 00 00 00 00
EG-N $-0.7$ $-1.3$ $-1.1$ $-1.3$ $0.3$ $-0.6$ $-2.0$ $7.4$ $7.3$ $7.4$ $7.3$ $7.4$ $7.4$ $7.3$ $1$ $100$ $100$
EG-T $0.5$ $-0.5$ $0.0$ $-0.5$ $2.3$ $2.0$ $-0.5$ $6.4$ $6.4$ $6.4$ $6.4$ $6.4$ $6.4$ $6.4$ $2.7$ $100$ $100$
REG-N $-5.4$ $-4.4$ $-5.2$ $-4.7$ $-6.4$ $-5.7$ $-6.2$ $9.1$ $9.2$ $9.0$ $9.3$ $8.2$ $8.6$ $8.9$ $-6.9$ $-6.0$ $100$
REG-T $-5.5$ $-4.5$ $-5.3$ $-4.7$ $-6.4$ $-5.8$ $-6.3$ $9.0$ $9.0$ $8.8$ $9.1$ $8.0$ $8.5$ $8.7$ $-7.0$ $-6.1$ $-8.4$
  • Notes: EG and REG refer to the EGARCH and REGARCH models, respectively. The lower triangular sections report the test statistics from the unconditional GW test; a negative (positive) value indicates that the row (column) model performs better than the column (row) model. Conversely, the upper triangular sections present the percentage of predicted losses, as defined in equation (62); a lower (higher) percentage suggests that the row (column) model outperforms the column (row) model. Dark gray and light gray shading indicate rejection of the GW test at the 1% and 5% significance levels, respectively.
Table 4 presents results from the GW tests. The lower triangular part reports unconditional test statistics: negative values favor the row model. The upper triangular part shows conditional win percentages, where larger values indicate better performance by the column model. Shaded cells denote significance at the 1% (dark gray) and 5% (light gray) levels.
For the DJIA, SV-FS-SN and SV-FS-ST outperform other SV variants and EGARCH models, while RSV and REGARCH models generally dominate SV and EGARCH. RSV-AZ-ST is significantly superior to even FS-type SV models, although RSV and REGARCH differ less markedly. Conditional GW tests further confirm the superiority of RSV models, particularly over SV and EGARCH.
For the N225, both unconditional and conditional GW tests support the outperformance of RSV and REGARCH over SV and EGARCH, with RSV models often exceeding REGARCH. FS-type distributions enhance forecast accuracy within the SV and RSV frameworks. Among RSV specifications, RSV-GH-ST lags significantly.
In summary, incorporating RV substantially improves forecast accuracy, with RSV and REGARCH models consistently outperforming SV and EGARCH counterparts. These results corroborate prior findings (Takahashi et al., 2023, 2024). The added flexibility of FS-type distributions further improves forecast accuracy, especially in SV models. The next section examines whether these improvements carry over to tail risk measures, including VaR and ES.
4.3.2 VaR and ES forecasts
Figure 10 illustrates the 1% VaR and ES forecasts for the DJIA and N225 indices, as generated by the RSV-FS-ST and REGARCH-T models. VaR violations—instances where realized returns fall below the forecasted VaR—are observed across both models, occurring during periods of market stress such as early 2018 and late 2019, as well as during more tranquil phases. For clarity, the forecasts from other models, which follow similar patterns with minor differences in scale, are omitted.
RSV-FS-ST for DJIA RSV-FS-ST for N225
Refer to caption Refer to caption
REGARCH-T for DJIA REGARCH-T for N225
Refer to caption Refer to caption
Figure 10: 1% VaR and ES forecasts of RSV-FS-ST and REGARCH-T models for DJIA and N225.
Table 5 presents the empirical violation rates ($\hat{\alpha}$), the average FZ0 loss values (as defined in Equation (57) of Patton et al. (2019)), and the corresponding $p$-values from the dynamic quantile (DQ) test and the MCS procedure. At the 1% level, all models tend to underestimate tail risk, as indicated by violation rates exceeding the nominal level. In contrast, for $\alpha=5\%$, the violation rates are generally well aligned with the target.
Regarding predictive accuracy based on the FZ0 loss, the SV-AZ-SN and RSV-AZ-ST models achieve the lowest losses at $\alpha=1\%$ and 5%, respectively. Most models—except SV-GH-ST at $\alpha=5\%$—are included in the 75% MCS, indicating comparable performance in tail risk forecasting.
The DQ test by Engle and Manganelli (2004), which regresses VaR violations on a constant, the forecasted VaR, and a one-period lagged violation indicator, suggests that the SV-FS-ST and SV-GH-ST models fail the independence test at the 1% level. This implies potential misspecification in their dynamic structures.
Table 5: Violation rates and FZ0 losses of VaR and ES forecasts for DJIA
$\alpha=1\%$ $\alpha=5\%$
$\hat{\alpha}$ $p_{DQ}$ FZ0 $p_{MCS}$ $\hat{\alpha}$ $p_{DQ}$ FZ0 $p_{MCS}$
SV-N $2.32$ $0.17$ $1.0714$ $1.00$ $4.64$ $0.97$ $0.5965$ $0.80$
SV-T $1.99$ $0.12$ $1.0483$ $1.00$ $4.31$ $0.89$ $0.5809$ $1.00$
SV-AZ-SN $1.82$ $0.08$ $0.9698$ $1.00$ $4.31$ $0.89$ $0.5858$ $0.99$
SV-AZ-ST $1.82$ $0.08$ $1.0180$ $1.00$ $4.31$ $0.65$ $0.5830$ $1.00$
SV-FS-SN $1.49$ $0.93$ $0.9786$ $1.00$ $4.15$ $0.84$ $0.5947$ $0.92$
SV-FS-ST $1.49$ $0.02$ $0.9899$ $1.00$ $4.15$ $0.84$ $0.5954$ $0.87$
SV-GH-ST $1.33$ $0.01$ $1.1224$ $1.00$ $4.31$ $0.92$ $0.7012$ $0.11$
RSV-N $2.82$ $0.23$ $1.2345$ $0.41$ $5.80$ $0.67$ $0.5567$ $1.00$
RSV-T $2.65$ $0.23$ $1.1709$ $0.82$ $5.80$ $0.57$ $0.5480$ $1.00$
RSV-AZ-SN $2.32$ $0.19$ $1.1275$ $1.00$ $5.47$ $0.67$ $0.5411$ $1.00$
RSV-AZ-ST $2.16$ $0.16$ $1.0780$ $1.00$ $5.14$ $0.68$ $0.5299$ $1.00$
RSV-FS-SN $2.32$ $0.19$ $1.0490$ $1.00$ $5.64$ $0.64$ $0.5411$ $1.00$
RSV-FS-ST $2.49$ $0.21$ $1.0958$ $1.00$ $5.31$ $0.63$ $0.5365$ $1.00$
RSV-GH-ST $1.99$ $0.12$ $1.0186$ $1.00$ $4.98$ $0.58$ $0.5307$ $1.00$
EGARCH-N $2.65$ $0.22$ $1.2928$ $0.40$ $4.81$ $0.90$ $0.6226$ $0.27$
EGARCH-T $1.82$ $0.83$ $1.0797$ $1.00$ $4.98$ $0.92$ $0.5855$ $1.00$
REGARCH-N $3.15$ $0.24$ $1.3273$ $0.25$ $5.47$ $0.81$ $0.5652$ $1.00$
REGARCH-T $2.99$ $0.23$ $1.3202$ $0.27$ $5.47$ $0.81$ $0.5621$ $1.00$
  • Notes: $\hat{\alpha}$ represents the empirical violation rate (%). $p_{DQ}$ indicates the $p$-value of the dynamic quantile test of Engle and Manganelli (2004). FZ0 denotes the average FZ0 loss. $p_{MCS}$ indicates the MCS $p$-value.
Table 6 summarizes the results for the N225 index. Across both target levels, most models yield violation rates consistent with the nominal levels. The RSV-AZ-ST model again performs best in terms of FZ0 loss, and both RSV and REGARCH models are included in the 75% MCS. Notably, the SV-GH-ST and EGARCH-N models fall outside the MCS at the 5% level, indicating weaker predictive accuracy. Unlike the DJIA case, none of the models are rejected by the DQ test for N225.
Table 6: Violation rates and FZ0 losses of VaR and ES forecasts for N225
$\alpha=1\%$ $\alpha=5\%$
$\hat{\alpha}$ $p_{DQ}$ FZ0 $p_{MCS}$ $\hat{\alpha}$ $p_{DQ}$ FZ0 $p_{MCS}$
SV-N $1.02$ $1.00$ $1.2034$ $0.90$ $3.90$ $0.85$ $0.8163$ $0.43$
SV-T $0.85$ $0.99$ $1.2158$ $0.55$ $3.39$ $0.64$ $0.7973$ $1.00$
SV-AZ-SN $0.85$ $0.99$ $1.2047$ $0.90$ $3.56$ $0.71$ $0.8178$ $0.51$
SV-AZ-ST $0.85$ $0.99$ $1.2053$ $0.95$ $3.56$ $0.75$ $0.8156$ $0.60$
SV-FS-SN $0.85$ $0.98$ $1.1979$ $1.00$ $3.39$ $0.65$ $0.8105$ $0.87$
SV-FS-ST $0.85$ $0.97$ $1.1870$ $1.00$ $3.39$ $0.64$ $0.8093$ $0.96$
SV-GH-ST $0.68$ $0.96$ $1.2720$ $0.27$ $4.24$ $0.87$ $0.8846$ $0.00$
RSV-N $1.53$ $0.83$ $1.1468$ $1.00$ $5.08$ $0.99$ $0.7854$ $1.00$
RSV-T $1.36$ $0.87$ $1.1357$ $1.00$ $4.92$ $0.98$ $0.7833$ $1.00$
RSV-AZ-SN $1.36$ $0.84$ $1.1544$ $1.00$ $4.75$ $0.97$ $0.7857$ $1.00$
RSV-AZ-ST $1.19$ $0.82$ $1.1260$ $1.00$ $4.75$ $0.97$ $0.7822$ $1.00$
RSV-FS-SN $1.19$ $0.83$ $1.1533$ $1.00$ $5.25$ $0.99$ $0.7887$ $1.00$
RSV-FS-ST $1.19$ $0.83$ $1.1301$ $1.00$ $5.25$ $1.00$ $0.7865$ $1.00$
RSV-GH-ST $1.19$ $0.83$ $1.1283$ $1.00$ $4.75$ $0.98$ $0.7871$ $1.00$
EGARCH-N $1.02$ $0.93$ $1.2818$ $0.58$ $3.73$ $0.79$ $0.8450$ $0.00$
EGARCH-T $1.02$ $0.91$ $1.2344$ $0.29$ $3.73$ $0.77$ $0.8292$ $0.21$
REGARCH-N $1.69$ $0.71$ $1.1844$ $1.00$ $4.41$ $0.94$ $0.7997$ $1.00$
REGARCH-T $1.02$ $0.79$ $1.1306$ $1.00$ $4.92$ $0.99$ $0.7959$ $1.00$
  • Notes: See Table 5 for additional details.
Figure 11 shows the CLDs relative to the SV-N benchmark. For the DJIA, REGARCH models perform poorly during the early 2018 volatility spike at the 1% level. From mid-2018 onwards, REGARCH and RSV models outperform SV and EGARCH models at the 5% level. For the N225, the superiority of RSV and REGARCH models is consistent throughout the forecast horizon, reinforcing the benefit of incorporating RV.
$\alpha=1\%$ for DJIA $\alpha=5\%$ for DJIA
Refer to caption Refer to caption
$\alpha=1\%$ for N225 $\alpha=5\%$ for N225
Refer to caption Refer to caption
Figure 11: Cumulative loss differences (FZ0) relative to the SV-N model for DJIA and N225.
GW test results
Table 7 reports the GW test results for the FZ0 loss function for the DJIA. The interpretation of the lower and upper triangular elements follows that of Table 4. The main findings are summarized below.
At the 1% level, the unconditional GW test reveals that SV-AZ-SN significantly outperforms SV-N, while SV-GH-ST is significantly outperformed by SV-AZ-SN. Among RSV models, RSV-T, RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, and RSV-GH-ST all show significantly better predictive accuracy than RSV-N. For EGARCH (EG) models, EG-T significantly outperforms EG-N. Regarding REGARCH (REG) models, both REG-N and REG-T are significantly outperformed by several RSV models, particularly RSV-AZ-SN, RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST.
The conditional GW test at the 1% level indicates that SV-T performs significantly worse than SV-AZ-SN, but better than EG-N. RSV-N is significantly outperformed by RSV-T, RSV-AZ-SN, RSV-FS-SN, and RSV-FS-ST. Moreover, all RSV variants (RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, RSV-GH-ST) significantly outperform REG-N and REG-T.
At the 5% level, the unconditional test shows that SV-GH-ST is significantly outperformed by SV-T, SV-AZ-SN, SV-AZ-ST, and all RSV models. In addition, RSV-GH-ST outperforms SV-N, and RSV-AZ-ST significantly outperforms RSV-AZ-SN. Among EGARCH models, EG-T again significantly outperforms EG-N. In the REGARCH group, REG-T outperforms REG-N, and both REGARCH variants outperform SV-GH-ST.
For the conditional GW test at the 5% level, SV-GH-ST is significantly outperformed by RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, and RSV-GH-ST. RSV-T performs significantly worse than RSV-GH-ST. EG-N is significantly outperformed by EG-T, and REG-N by REG-T.
In summary, the results highlight the superior performance of the RSV models—especially RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST—compared to SV, EGARCH, and REGARCH models, particularly under stringent risk levels. The inclusion of RV consistently improves the accuracy of tail risk forecasts. Moreover, as the FZ0 loss function jointly evaluates VaR and ES by emphasizing tail behavior, these results underscore the importance of capturing skewness and heavy tails in return distributions.
Table 7: Summary of GW tests on FZ0 loss of VaR and ES forecasts for DJIA
$\alpha=1\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $94$ $99$ $99$ $100$ $99$ $6$ $1$ $1$ $9$ $54$ $99$ $12$ $99$ $1$ $23$ $1$ $1$
SV-T $-0.4$ $99$ $100$ $35$ $100$ $7$ 00 00 00 $1$ $14$ $1$ $99$ $2$ $2$ 00 00
SV-AZ-SN $-2.2$ $-1.6$ $4$ $6$ 00 $2$ 00 00 00 $1$ 00 00 $1$ $1$ $1$ 00 00
SV-AZ-ST $-0.9$ $-1.0$ $1.4$ $47$ $59$ $4$ 00 00 00 00 $1$ 00 $12$ $2$ $1$ 00 00
SV-FS-SN $-1.5$ $-0.8$ $0.2$ $-0.6$ $11$ $2$ 00 00 $5$ $16$ 00 $4$ $49$ 00 00 00 00
SV-FS-ST $-1.4$ $-0.9$ $0.9$ $-0.7$ $0.3$ $2$ 00 00 00 00 00 00 $4$ $1$ $1$ 00 00
SV-GH-ST $0.5$ $0.8$ $2.0$ $1.5$ $1.5$ $1.9$ $2$ $2$ $4$ $89$ $96$ $57$ $97$ $2$ $71$ 00 00
RSV-N $1.1$ $1.4$ $1.6$ $1.5$ $1.4$ $1.4$ $0.6$ $100$ $100$ $100$ $100$ $100$ $100$ $2$ $99$ $2$ $2$
RSV-T $0.8$ $1.1$ $1.4$ $1.2$ $1.2$ $1.2$ $0.3$ $-2.2$ $100$ $99$ $100$ $100$ $100$ $2$ $99$ 00 00
RSV-AZ-SN $0.4$ $0.8$ $1.2$ $0.9$ $1.0$ $1.0$ $0.0$ $-2.3$ $-1.6$ $100$ $97$ $97$ $100$ $2$ $99$ 00 00
RSV-AZ-ST $0.1$ $0.3$ $1.0$ $0.6$ $0.8$ $0.7$ $-0.3$ $-2.3$ $-2.0$ $-1.7$ $90$ 00 $100$ $2$ $62$ 00 00
RSV-FS-SN $-0.2$ $0.0$ $0.9$ $0.3$ $0.7$ $0.6$ $-0.6$ $-2.1$ $-1.8$ $-1.2$ $-0.6$ $1$ $100$ $2$ $6$ 00 00
RSV-FS-ST $0.2$ $0.5$ $1.1$ $0.8$ $0.9$ $0.8$ $-0.2$ $-2.4$ $-2.1$ $-1.2$ $0.9$ $1.1$ $100$ $2$ $98$ 00 00
RSV-GH-ST $-0.6$ $-0.4$ $0.6$ $0.0$ $0.4$ $0.3$ $-0.9$ $-2.2$ $-2.0$ $-1.8$ $-1.7$ $-0.8$ $-1.8$ $2$ $1$ 00 00
EG-N $1.5$ $1.8$ $1.9$ $1.8$ $1.6$ $1.7$ $0.9$ $0.4$ $0.8$ $1.1$ $1.3$ $1.5$ $1.3$ $1.7$ $100$ $1$ $1$
EG-T $0.1$ $0.4$ $1.2$ $0.7$ $0.9$ $0.9$ $-0.3$ $-1.0$ $-0.6$ $-0.3$ $0.0$ $0.3$ $-0.1$ $0.5$ $-2.1$ 00 00
REG-N $1.4$ $1.7$ $1.8$ $1.7$ $1.5$ $1.6$ $0.9$ $1.3$ $1.8$ $2.2$ $2.2$ $2.0$ $2.1$ $2.2$ $0.2$ $1.4$ $98$
REG-T $1.4$ $1.6$ $1.7$ $1.6$ $1.5$ $1.5$ $0.9$ $1.3$ $1.8$ $2.1$ $2.2$ $1.9$ $2.1$ $2.2$ $0.2$ $1.4$ $-1.8$
$\alpha=5\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $96$ $99$ $98$ $95$ $92$ $1$ $97$ $97$ $98$ $98$ $99$ $98$ $98$ $2$ $80$ $95$ $96$
SV-T $-0.6$ $7$ $5$ $7$ $5$ 00 $99$ $99$ $99$ $99$ $99$ $99$ $99$ $2$ $25$ $98$ $98$
SV-AZ-SN $-0.8$ $0.2$ $86$ $4$ 00 00 $97$ $98$ $98$ $98$ $100$ $99$ $99$ $2$ $38$ $96$ $96$
SV-AZ-ST $-0.6$ $0.2$ $-0.2$ $8$ $4$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $3$ $32$ $97$ $97$
SV-FS-SN $-0.1$ $0.5$ $0.8$ $0.5$ $34$ $1$ $96$ $97$ $97$ $98$ $98$ $98$ $98$ 00 $63$ $95$ $96$
SV-FS-ST $-0.1$ $0.6$ $0.9$ $0.7$ $0.1$ 00 $98$ $98$ $98$ $98$ $100$ $100$ $98$ $3$ $53$ $96$ $96$
SV-GH-ST $1.7$ $2.0$ $2.0$ $2.0$ $1.7$ $1.8$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $98$ $98$ $100$ $100$
RSV-N $-1.0$ $-0.6$ $-0.7$ $-0.7$ $-0.9$ $-0.9$ $-2.1$ $99$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $11$ $14$
RSV-T $-1.3$ $-0.9$ $-1.0$ $-0.9$ $-1.1$ $-1.1$ $-2.3$ $-1.7$ $96$ $96$ $82$ $100$ $98$ $1$ $1$ $5$ $6$
RSV-AZ-SN $-1.6$ $-1.3$ $-1.3$ $-1.3$ $-1.4$ $-1.4$ $-2.4$ $-1.4$ $-0.8$ $100$ $73$ $80$ $100$ $2$ $1$ $2$ $3$
RSV-AZ-ST $-1.9$ $-1.6$ $-1.6$ $-1.7$ $-1.7$ $-1.7$ $-2.6$ $-1.8$ $-1.5$ $-2.2$ $1$ $12$ $15$ $2$ $2$ 00 $1$
RSV-FS-SN $-1.7$ $-1.2$ $-1.3$ $-1.3$ $-1.5$ $-1.4$ $-2.4$ $-1.5$ $-0.7$ $0.0$ $1.1$ $100$ $99$ $1$ $2$ $8$ $9$
RSV-FS-ST $-1.8$ $-1.4$ $-1.5$ $-1.5$ $-1.6$ $-1.6$ $-2.5$ $-1.8$ $-1.2$ $-0.7$ $0.8$ $-0.8$ $100$ $2$ $2$ $5$ $7$
RSV-GH-ST $-2.0$ $-1.8$ $-1.8$ $-1.8$ $-1.9$ $-1.9$ $-2.7$ $-1.4$ $-1.1$ $-1.2$ $0.1$ $-0.9$ $-0.6$ $2$ $2$ $1$ $1$
EG-N $0.8$ $1.5$ $1.1$ $1.3$ $0.8$ $0.8$ $-1.1$ $1.4$ $1.6$ $1.9$ $2.1$ $1.9$ $2.0$ $2.2$ $99$ $99$ $99$
EG-T $-0.4$ $0.2$ $0.0$ $0.1$ $-0.4$ $-0.4$ $-1.7$ $0.6$ $0.8$ $1.0$ $1.2$ $1.0$ $1.1$ $1.3$ $-2.3$ $99$ $100$
REG-N $-0.8$ $-0.5$ $-0.5$ $-0.5$ $-0.7$ $-0.7$ $-2.0$ $0.5$ $1.2$ $1.7$ $2.2$ $1.2$ $1.6$ $1.8$ $-1.4$ $-0.4$ $97$
REG-T $-0.8$ $-0.5$ $-0.6$ $-0.6$ $-0.7$ $-0.7$ $-2.0$ $0.3$ $1.0$ $1.5$ $2.1$ $1.1$ $1.5$ $1.7$ $-1.5$ $-0.5$ $-2.5$
  • Notes: See Table 4 for additional details.
Table 8 summarizes the GW test results for the N225, which are broadly consistent with those for the DJIA. At the 1% level, the unconditional test shows that several RSV models significantly outperform SV models, particularly SV-GH-ST, while REGARCH models exhibit mixed performance relative to RSV models. The conditional GW test confirms that RSV models consistently outperform both SV-GH-ST and REGARCH models.
At the 5% level, the unconditional test results indicate that SV-GH-ST is significantly outperformed by all SV models and all RSV models. Although REGARCH models perform better than SV-GH-ST, they are still outperformed by RSV models. The conditional test further supports the dominance of RSV models over SV, EGARCH, and REGARCH variants.
Taken together, the N225 results reinforce the conclusions drawn from the DJIA, providing additional evidence that RSV models—especially RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST—consistently deliver superior VaR and ES forecasts across varying levels of tail risk.
Table 8: Summary of GW tests on FZ0 loss of VaR and ES forecasts for N225
$\alpha=1\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $1$ $1$ 00 $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ 00 $100$ $100$
SV-T $0.7$ $100$ $99$ $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ 00 $100$ $100$
SV-AZ-SN $0.1$ $-0.5$ $1$ $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $100$ $99$
SV-AZ-ST $0.1$ $-0.4$ $0.0$ $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $100$ $99$
SV-FS-SN $-0.1$ $-0.5$ $-0.3$ $-0.3$ $100$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $100$ $99$
SV-FS-ST $-0.3$ $-0.8$ $-0.5$ $-0.8$ $-0.4$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $77$ $99$
SV-GH-ST $1.1$ $1.1$ $1.3$ $1.8$ $1.9$ $2.3$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $94$ $99$ $99$
RSV-N $-1.0$ $-1.1$ $-0.9$ $-0.8$ $-0.7$ $-0.5$ $-1.5$ $99$ 00 $99$ 00 $99$ $99$ 00 00 00 $99$
RSV-T $-1.4$ $-1.5$ $-1.4$ $-1.2$ $-0.9$ $-0.7$ $-1.8$ $-0.5$ 00 $100$ 00 $100$ $100$ 00 00 $1$ $99$
RSV-AZ-SN $-0.8$ $-1.0$ $-0.8$ $-0.7$ $-0.6$ $-0.4$ $-1.4$ $1.0$ $0.7$ $99$ $99$ $99$ $99$ 00 00 $1$ $99$
RSV-AZ-ST $-1.7$ $-2.0$ $-1.7$ $-1.6$ $-1.2$ $-1.0$ $-2.2$ $-0.9$ $-0.5$ $-1.2$ 00 $1$ 00 $1$ 00 $1$ 00
RSV-FS-SN $-1.2$ $-1.4$ $-1.1$ $-1.0$ $-0.8$ $-0.5$ $-1.7$ $0.3$ $1.1$ $0.0$ $1.9$ $100$ $99$ $1$ 00 $2$ $99$
RSV-FS-ST $-1.5$ $-1.7$ $-1.6$ $-1.4$ $-1.1$ $-0.9$ $-2.1$ $-0.8$ $-0.4$ $-1.1$ $0.3$ $-1.7$ $99$ 00 00 $1$ $72$
RSV-GH-ST $-1.6$ $-1.9$ $-1.7$ $-1.6$ $-1.2$ $-1.0$ $-2.2$ $-0.7$ $-0.4$ $-0.9$ $0.3$ $-1.4$ $-0.2$ $1$ 00 $1$ $3$
EG-N $1.0$ $0.8$ $0.8$ $0.8$ $0.7$ $0.9$ $0.1$ $1.7$ $1.8$ $1.6$ $1.8$ $1.6$ $1.7$ $1.7$ $99$ $100$ $99$
EG-T $0.9$ $0.8$ $0.8$ $0.8$ $0.7$ $1.2$ $-0.7$ $1.4$ $1.9$ $1.3$ $2.2$ $1.7$ $2.0$ $2.1$ $-0.7$ $100$ $100$
REG-N $-0.3$ $-0.5$ $-0.3$ $-0.3$ $-0.2$ $0.0$ $-1.0$ $2.2$ $1.5$ $1.8$ $1.7$ $1.0$ $1.7$ $1.5$ $-1.4$ $-0.8$ $99$
REG-T $-1.7$ $-2.2$ $-2.0$ $-2.1$ $-1.6$ $-1.2$ $-2.7$ $-0.4$ $-0.2$ $-0.6$ $0.2$ $-0.9$ $0.0$ $0.1$ $-1.6$ $-2.4$ $-1.0$
$\alpha=5\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $99$ $5$ $53$ $97$ $100$ $1$ $99$ $100$ $99$ $100$ $100$ $99$ $99$ $1$ $2$ $98$ $100$
SV-T $-2.0$ 00 $1$ $1$ $3$ 00 $100$ $100$ $99$ $100$ $99$ $100$ $98$ $1$ $2$ $15$ $99$
SV-AZ-SN $0.2$ $3.2$ $97$ $97$ $99$ $1$ $98$ $99$ $98$ $98$ $100$ $98$ $98$ $2$ $2$ $97$ $98$
SV-AZ-ST $-0.1$ $3.2$ $-0.3$ $99$ $86$ 00 $99$ $100$ $99$ $100$ $100$ $99$ $99$ $1$ $2$ $97$ $100$
SV-FS-SN $-0.6$ $1.4$ $-1.1$ $-0.7$ $99$ $1$ $97$ $99$ $97$ $97$ $99$ $97$ $97$ $1$ $1$ $96$ $97$
SV-FS-ST $-0.5$ $1.3$ $-0.8$ $-0.8$ $-0.1$ $1$ $97$ $97$ $97$ $97$ $97$ $97$ $97$ $1$ $6$ $94$ $96$
SV-GH-ST $2.2$ $3.0$ $2.2$ $2.4$ $2.4$ $2.5$ $99$ $99$ $99$ $99$ $99$ $99$ $99$ $96$ $98$ $99$ $99$
RSV-N $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.6$ $99$ $3$ $100$ $15$ $3$ 00 00 00 00 00
RSV-T $-1.3$ $-0.5$ $-1.3$ $-1.2$ $-1.0$ $-0.8$ $-2.7$ $-0.6$ $3$ $98$ 00 $2$ $1$ 00 00 $1$ 00
RSV-AZ-SN $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.6$ $0.1$ $0.7$ $99$ $1$ $2$ 00 00 00 00 00
RSV-AZ-ST $-1.4$ $-0.6$ $-1.3$ $-1.2$ $-1.0$ $-0.9$ $-2.8$ $-0.7$ $-0.3$ $-0.8$ 00 $1$ 00 00 00 00 00
RSV-FS-SN $-1.1$ $-0.3$ $-1.0$ $-0.9$ $-0.7$ $-0.6$ $-2.6$ $0.8$ $1.1$ $0.8$ $1.1$ $99$ $100$ 00 00 $2$ 00
RSV-FS-ST $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.7$ $0.3$ $1.1$ $0.2$ $1.0$ $-0.4$ $38$ 00 00 00 00
RSV-GH-ST $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.7$ $0.4$ $1.0$ $0.3$ $1.9$ $-0.3$ $0.1$ 00 00 00 $1$
EG-N $1.2$ $2.2$ $1.1$ $1.2$ $1.3$ $1.4$ $-1.3$ $2.0$ $2.2$ $2.1$ $2.2$ $2.0$ $2.1$ $2.1$ $100$ $100$ $100$
EG-T $0.6$ $2.0$ $0.6$ $0.8$ $0.9$ $1.1$ $-2.0$ $1.5$ $1.7$ $1.5$ $1.7$ $1.4$ $1.5$ $1.6$ $-1.5$ $100$ $100$
REG-N $-0.7$ $0.1$ $-0.7$ $-0.6$ $-0.4$ $-0.3$ $-2.4$ $2.0$ $2.7$ $2.1$ $3.0$ $1.3$ $2.1$ $2.6$ $-1.8$ $-1.2$ $98$
REG-T $-0.9$ $-0.1$ $-0.9$ $-0.8$ $-0.6$ $-0.5$ $-2.6$ $1.2$ $1.9$ $1.3$ $2.4$ $0.8$ $1.4$ $1.7$ $-1.8$ $-1.4$ $-0.8$
  • Notes: See Table 4 for additional details.
Distributional characteristics of daily return forecasts under RSV models
To explore differences among RSV models, we examine the distributional properties of their one-day-ahead return forecasts. Table 9 reports the mean and standard deviation of skewness, kurtosis, and several lower percentiles (0.1%, 1%, 5%, and 10%) based on the posterior predictive distributions. Figures 12 and 13 present the histograms of these statistics across the forecast horizon for the DJIA and N225, respectively.
For both indices, the RSV-FS-SN and RSV-GH-ST models tend to generate more negatively skewed and heavy-tailed predictive distributions, as reflected in their higher kurtosis and lower percentile values. Among all models, RSV-GH-ST exhibits the most negative skewness, while RSV-FS-SN displays the highest kurtosis.
For the N225 index, the histograms of the 0.1st and 1st percentiles indicate that RSV-FS-SN consistently produces more conservative forecasts in the extreme left tail. Interestingly, its 10th percentile is relatively higher than those of other models. These distributional characteristics may help explain the variation in VaR and ES forecast performance observed across RSV models.
Table 9: Mean characteristics of one-day-ahead return forecasts under RSV models
DJIA
Skewness Kurtosis $0.1\%$ $1\%$ $5\%$ $10\%$
RSV-N $0.000$ $(0.03)$ $3.588$ $(0.36)$ $-3.473$ $(0.18)$ $-2.436$ $(0.05)$ $-1.633$ $(0.02)$ $-1.245$ $(0.02)$
RSV-T $0.002$ $(0.04)$ $4.050$ $(0.20)$ $-3.716$ $(0.15)$ $-2.496$ $(0.03)$ $-1.624$ $(0.01)$ $-1.222$ $(0.01)$
RSV-AZ-SN $-0.241$ $(0.05)$ $3.691$ $(0.10)$ $-3.783$ $(0.13)$ $-2.599$ $(0.04)$ $-1.695$ $(0.02)$ $-1.265$ $(0.01)$
RSV-AZ-ST $-0.255$ $(0.06)$ $4.166$ $(0.17)$ $-4.019$ $(0.16)$ $-2.656$ $(0.04)$ $-1.684$ $(0.02)$ $-1.242$ $(0.01)$
RSV-FS-SN $-0.244$ $(0.07)$ $4.764$ $(0.35)$ $-4.225$ $(0.19)$ $-2.690$ $(0.05)$ $-1.659$ $(0.02)$ $-1.219$ $(0.01)$
RSV-FS-ST $-0.216$ $(0.05)$ $4.004$ $(0.22)$ $-3.894$ $(0.17)$ $-2.613$ $(0.04)$ $-1.679$ $(0.02)$ $-1.249$ $(0.01)$
RSV-GH-ST $-0.388$ $(0.07)$ $4.390$ $(0.32)$ $-4.250$ $(0.20)$ $-2.716$ $(0.05)$ $-1.694$ $(0.02)$ $-1.241$ $(0.01)$
N225
Skewness Kurtosis $0.1\%$ $1\%$ $5\%$ $10\%$
RSV-N $-0.001$ $(0.03)$ $3.487$ $(0.08)$ $-3.420$ $(0.10)$ $-2.425$ $(0.03)$ $-1.636$ $(0.01)$ $-1.249$ $(0.01)$
RSV-T $0.001$ $(0.04)$ $3.892$ $(0.14)$ $-3.639$ $(0.12)$ $-2.475$ $(0.03)$ $-1.627$ $(0.01)$ $-1.229$ $(0.01)$
RSV-AZ-SN $-0.017$ $(0.03)$ $3.498$ $(0.08)$ $-3.451$ $(0.10)$ $-2.435$ $(0.03)$ $-1.639$ $(0.01)$ $-1.251$ $(0.01)$
RSV-AZ-ST $-0.019$ $(0.03)$ $3.896$ $(0.13)$ $-3.665$ $(0.12)$ $-2.490$ $(0.04)$ $-1.632$ $(0.01)$ $-1.231$ $(0.01)$
RSV-FS-SN $-0.064$ $(0.07)$ $4.661$ $(0.28)$ $-4.041$ $(0.20)$ $-2.577$ $(0.05)$ $-1.624$ $(0.02)$ $-1.204$ $(0.01)$
RSV-FS-ST $-0.066$ $(0.05)$ $3.850$ $(0.17)$ $-3.670$ $(0.16)$ $-2.510$ $(0.04)$ $-1.645$ $(0.02)$ $-1.238$ $(0.01)$
RSV-GH-ST $-0.081$ $(0.04)$ $3.903$ $(0.15)$ $-3.749$ $(0.14)$ $-2.524$ $(0.04)$ $-1.644$ $(0.01)$ $-1.235$ $(0.01)$
  • Notes: The table reports the means of skewness, kurtosis, and lower percentiles ($0.1\%$, $1\%$, $5\%$, and $10\%$) computed from the posterior predictive distributions of one-day-ahead return forecasts. Standard deviations are shown in parentheses. These statistics are calculated across 603 and 590 forecast days for the DJIA and N225, respectively.
Figure 12: Histograms of skewness, kurtosis, and lower percentiles ($0.1\%$, $1\%$, $5\%$, and $10\%$) computed from the posterior predictive distributions of one-day-ahead return forecasts for the DJIA.
Figure 13: Histograms of skewness, kurtosis, and lower percentiles ($0.1\%$, $1\%$, $5\%$, and $10\%$) computed from the posterior predictive distributions of one-day-ahead return forecasts for the N225.
4.3.3 Summary
Table 10 presents the rankings of average losses for volatility, VaR, and ES forecasts. The average ranking in the final column highlights the overall superiority of the RSV models, with RSV-AZ-ST achieving the best performance.
Our findings demonstrate that RV is a valuable predictor for forecasting not only volatility but also tail risk measures such as VaR and ES. Across all forecast categories, RSV models consistently outperform REGARCH models. In particular, incorporating a skewed $t$-distribution into the RSV framework appears to further enhance forecast accuracy.
Table 10: Ranking of average losses for volatility, VaR, and ES forecasts
DJIA N225
QLIKE FZ0 QLIKE FZ0
RV5 RK BV Med 1% 5% RV5 RK BV Med 1% 5% Average
SV-N 17 17 17 15   8 16 13 12 13 14 12 14 $14.00$
SV-T 14 14 14 14   6 10 18 18 16 15 15   9 $13.58$
SV-AZ-SN 18 18 18 18   1 13 15 14 14 16 13 15 $14.42$
SV-AZ-ST 15 15 15 16   4 11 17 16 17 17 14 13 $14.17$
SV-FS-SN 11   2   2   2   2 14 10 10 10 11 11 12 $8.08$
SV-FS-ST 10   1   1   1   3 15 12 13 12 12 10 11 $8.42$
SV-GH-ST 16 16 16 17 12 18 16 15 18 18 17 18 $16.42$
RSV-N   6 12 10 11 15   7   3   3   3   5   6   3 $7.00$
RSV-T   3   7   5   5 14   6   5   5   5   3   5   2 $5.42$
RSV-AZ-SN   4 10   8   9 13   4   4   4   4   4   8   4 $6.33$
RSV-AZ-ST   2   5   4   4   9   1   6   6   6   1   1   1 $3.83$
RSV-FS-SN   9 13 12 13   7   5   1   1   1   7   7   7 $6.92$
RSV-FS-ST   5 11   9 10 11   3   2   2   2   6   3   5 $5.75$
RSV-GH-ST   1   3   3   3   5   2   7   7   7   2   2   6 $4.00$
EGARCH-N 13   9 13 12 16 17 11 11 11 10 18 17 $13.17$
EGARCH-T 12   8 11   8 10 12 14 17 15 13 16 16 $12.67$
REGARCH-N   8   6   7   7 18   9   9   9   9   9   9 10 $9.17$
REGARCH-T   7   4   6   6 17   8   8   8   8   8   4   8 $7.67$
  • Notes: QLIKE denotes the Gaussian Quasi Likelihood loss for volatility forecasts using four volatility proxies: RV5, RK, BV, and Med. FZ0 refers to the FZ0 loss functions for VaR and ES forecasts at significance levels $\alpha=1\%$ and $5\%$. The last column indicates the average rank across all categories.
5 Conclusion
This study has evaluated the predictive performance of multiple models for forecasting volatility, VaR, and ES using data from two major financial indices: the DJIA and N225. Our comprehensive assessment, summarized in Table 10, demonstrates that the RSV models consistently outperform alternative approaches across all forecast objectives.
Among the models analyzed, the RSV-AZ-ST specification achieves the best overall performance. This highlights not only the value of incorporating RV but also the benefit of accounting for skewness and heavy tails in return distributions. Specifically, the use of skewed $t$-distributions within the RSV framework substantially improves forecast accuracy for both volatility and tail risk measures.
A key contribution of this study is the joint validation of these modeling enhancements: integrating RV and adopting flexible return distributions. These improvements offer methodological guidance for constructing robust models for financial risk forecasting.
In sum, our findings suggest that RSV models—particularly those incorporating skewed $t$-distributions—constitute a robust and effective approach for forecasting volatility, VaR, and ES. These results emphasize the critical role of model specification and distributional assumptions in improving the reliability of financial risk forecasts.
Appendix
Appendix A MCMC Sampling Scheme
A.1 RSV-AZ-ST Model
The joint probability density of $(y_{t},h_{t+1})$ conditional on $z_{0t}$, $h_{t}$, and parameters $\bm{\theta}$ is given by
$f(y_{t},h_{t+1}|z_{0t},h_{t},\bm{\theta})=$ $\frac{1}{2\pi k(\delta)\sqrt{{(1-\rho^{2})\sigma_{\eta}^{2}}}}\sqrt{\frac{\mu_{\lambda}}{\lambda_{t}}}\exp\left(-\frac{1}{2}h_{t}\right)$
×$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\hskip 128.0374pt\left.\left.-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\rho\exp(h_{t}/2)\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]^{2}\right\}$
×$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}\right\},$
where
$k(\delta)=\sqrt{\frac{1-\delta^{2}}{1-c^{2}\delta^{2}}}.$
The full conditional posterior density is then given by
$\pi(\bm{h},\bm{\theta},\bm{z}_{0},\bm{\lambda}|\bm{x},\bm{y})\propto$ $k(\delta)^{-T}(1-\rho^{2})^{-\frac{T-1}{2}}\sigma_{\eta}^{-T}(1-\phi^{2})^{\frac{1}{2}}$
×$\times\exp\left\{-\frac{1}{2}\sum_{t=1}^{T}h_{t}-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\hskip 136.5733pt\left.\left.-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\rho\exp(h_{t}/2)\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]^{2}\right\}$
×$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{(1-\phi^{2})}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
×$\times\exp\left\{-\frac{1}{2}\sum_{t=1}^{T}z_{0t}^{2}I(z_{0t}>0)\right\}\times\prod_{t=1}^{T}\frac{\mu_{\lambda}^{1/2}(\nu/2)^{\nu/2}}{\Gamma(\nu/2)}\lambda_{t}^{-(\nu+1)/2-1}\exp\left(-\frac{\nu}{2\lambda_{t}}\right)$
×$\times\sigma_{u}^{-T}\times\exp\left\{-\frac{1}{2\sigma_{u}^{2}}\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}\right\}\times\pi(\bm{\theta}),$
where $\pi(\bm{\theta})$ denotes the prior distribution of the model parameters.
A.1.1 Generation of $\mu$
The conditional posterior distribution of $\mu$ follows a normal distribution:
$\mu\mid\cdot\sim\mathcal{N}(\mu_{1},\sigma_{\mu 1}^{2}),$
where
$\sigma_{\mu 1}^{-2}$ $=\sigma_{\mu 0}^{-2}+\sigma_{\eta}^{-2}\left[1-\phi^{2}+(T-1)(1-\phi)^{2}(1-\rho^{2})^{-1}\right],$
$\mu_{1}$ $=\sigma_{\mu 1}^{2}\left\{\sigma_{\mu 0}^{-2}\mu_{0}+\sigma_{\eta}^{-2}\left[(1-\phi^{2})h_{1}+\frac{1-\phi}{1-\rho^{2}}\sum_{t=1}^{T-1}\left(h_{t+1}-\phi h_{t}\right.\right.\right.$
$\hskip 122.34692pt\left.\left.\left.-\sqrt{\frac{\mu_{\lambda}}{\lambda_{t}}}\cdot\frac{\rho\sigma_{\eta}\exp(-h_{t}/2)}{k(\delta)}\cdot\left(y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right)\right)\right]\right\}.$
A.1.2 Generation of $\phi$
The conditional posterior density of $\phi$ is given by
$\pi(\phi\mid\cdot)\propto$ $(1+\phi)^{a_{\phi 0}-1}(1-\phi)^{b_{\phi 0}-1}(1-\phi^{2})^{1/2}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\hskip 142.26378pt\left.\left.-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\rho\cdot\exp(h_{t}/2)\cdot\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]^{2}\right\}$
×$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
[…]
where
$g(\phi)$ $=(1+\phi)^{a_{\phi 0}-1/2}(1-\phi)^{b_{\phi 0}-1/2},$
$\sigma_{\phi}^{2}$ $=\frac{\sigma_{\eta}^{2}(1-\rho^{2})}{\rho^{2}(h_{1}-\mu)^{2}+\sum_{t=2}^{T-1}(h_{t}-\mu)^{2}},$
$\mu_{\phi}$ $=\frac{\sigma_{\phi}^{2}}{\sigma_{\eta}^{2}(1-\rho^{2})}\sum_{t=1}^{T-1}\left[(h_{t+1}-\mu)-\sqrt{\frac{\mu_{\lambda}}{\lambda_{t}}}\cdot\frac{\rho\sigma_{\eta}\exp(-h_{t}/2)}{k(\delta)}\left(y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right)\right](h_{t}-\mu).$
We then propose a candidate $\phi^{\dagger}\sim\mathcal{N}(\mu_{\phi},\sigma_{\phi}^{2})$ and accept it with probability $\min\left\{1,\frac{g(\phi^{\dagger})}{g(\phi)}\right\}$.
Generation of $\sigma_{\eta}^{2}$ and $\rho$
Instead of sampling $\sigma_{\eta}^{2}$ and $\rho$ separately, we sample them jointly. Let $\vartheta=(\sigma_{\eta},\rho)^{\prime}$. The conditional posterior density of $\vartheta$ is given by
$\pi(\vartheta\mid\cdot)$ $\propto(\sigma_{\eta}^{2})^{-(n_{0}+T)/2-1}\exp\left(-\frac{S_{0}}{2\sigma_{\eta}^{2}}\right)(1+\rho)^{a_{\rho 0}-1}(1-\rho)^{b_{\rho 0}-1}(1-\rho^{2})^{-(T-1)/2}$
×$\quad\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\hskip 153.6447pt\left.\left.-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\rho\cdot\exp(h_{t}/2)\cdot\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]^{2}\right\}$
×$\quad\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}.$
To handle the parameter constraints $\sigma_{\eta}>0$ and $|\rho|<1$, we perform a transformation:
$\omega_{1}=\log\sigma_{\eta},\quad\omega_{2}=\log(1+\rho)-\log(1-\rho).$
This transformation allows unconstrained sampling of $\omega=(\omega_{1},\omega_{2})$. Let $\hat{\vartheta}$ be the mode of the conditional posterior, and $\hat{\omega}$ its corresponding transformed value. We approximate the conditional posterior of $\omega$ by a normal distribution $\mathcal{N}(\omega_{*},\Sigma_{\eta*})$, where
$\omega_{*}$ $=\hat{\omega}+\Sigma_{\eta*}\cdot\left.\frac{\partial\log\pi(\omega\mid\cdot)}{\partial\omega}\right|_{\omega=\hat{\omega}},$
$\Sigma_{\eta*}^{-1}$ $=-\left.\frac{\partial^{2}\log\pi(\omega\mid\cdot)}{\partial\omega\partial\omega^{\prime}}\right|_{\omega=\hat{\omega}}.$
We then propose a candidate $\omega^{\dagger}\sim\mathcal{N}(\omega_{*},\Sigma_{\eta*})$, and accept it with probability
$\min\left\{1,\frac{\pi(\vartheta^{\dagger}\mid\cdot)\,f_{N}(\omega\mid\omega_{*},\Sigma_{\eta*})\,|J(\vartheta)|}{\pi(\vartheta\mid\cdot)\,f_{N}(\omega^{\dagger}\mid\omega_{*},\Sigma_{\eta*})\,|J(\vartheta^{\dagger})|}\right\},$
where $f_{N}(\cdot\mid\mu,\Sigma)$ denotes the density of a multivariate normal distribution, $J(\cdot)$ is the Jacobian of the inverse transformation, and $\vartheta^{\dagger}$ is the back-transformed candidate.
A.1.3 Generation of $\delta$
The conditional posterior density of $\delta$ is given by
$\pi(\delta\mid\cdot)\propto$ $(1+\delta)^{a_{\delta 0}-1}(1-\delta)^{b_{\delta 0}-1}\cdot k(\delta)^{-T}$
×$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\hskip 150.79968pt\left.\left.-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\rho\cdot\exp(h_{t}/2)\cdot\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]^{2}\right\}.$
We approximate the conditional posterior of $\delta$ by a normal distribution $\mathcal{N}(\hat{\delta},\sigma_{\hat{\delta}}^{2})$, where $\hat{\delta}$ is the mode of the conditional density and
$\sigma_{\hat{\delta}}^{-2}=-\left.\frac{\partial^{2}\log\pi(\delta\mid\cdot)}{\partial\delta^{2}}\right|_{\delta=\hat{\delta}}.$
We then propose a candidate value $\delta^{\dagger}\sim\mathcal{TN}_{(-1,1)}(\hat{\delta},\sigma_{\hat{\delta}}^{2})$ and accept it with probability
$\min\left\{1,\frac{\pi(\delta^{\dagger}\mid\cdot)\,f_{N}(\delta\mid\hat{\delta},\sigma_{\hat{\delta}}^{2})}{\pi(\delta\mid\cdot)\,f_{N}(\delta^{\dagger}\mid\hat{\delta},\sigma_{\hat{\delta}}^{2})}\right\},$
where $\mathcal{TN}_{(a,b)}(\mu,\sigma^{2})$ denotes a normal distribution truncated over the interval $(a,b)$, and $f_{N}(\cdot\mid\mu,\sigma^{2})$ is the density function of a normal distribution with mean $\mu$ and variance $\sigma^{2}$.
A.1.4 Generation of $\xi$
The conditional posterior distribution of $\xi$ is normal:
$\xi\mid\cdot\sim\mathcal{N}(\mu_{\xi},\sigma_{\xi}^{2}),$
where
$\sigma_{\xi}^{-2}$ $=\sigma_{\xi 0}^{-2}+T\sigma_{u}^{-2},$
$\mu_{\xi}$ $=\sigma_{\xi}^{2}\left[\sigma_{\xi 0}^{-2}\mu_{\xi 0}+\sigma_{u}^{-2}\sum_{t=1}^{T}(x_{t}-h_{t})\right].$
A.1.5 Generation of $\sigma_{u}^{2}$
The conditional posterior distribution of $\sigma_{u}^{2}$ follows an inverse gamma distribution:
$\sigma_{u}^{2}\mid\cdot\sim\mathcal{IG}\left(\frac{n_{u1}}{2},\frac{S_{u1}}{2}\right),$
where
$n_{u1}$ $=n_{u0}+T,$
$S_{u1}$ $=S_{u0}+\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}.$
A.1.6 Generation of $\nu$
The conditional posterior density of $\nu$ is given by
$\pi(\nu\mid\cdot)\propto$ $\exp\left\{\frac{T}{2}\log\mu_{\lambda}+\frac{\nu T}{2}\log\left(\frac{\nu}{2}\right)-T\log\Gamma\left(\frac{\nu}{2}\right)-\frac{\nu}{2}\sum_{t=1}^{T}\left(\lambda_{t}^{-1}+\log\lambda_{t}\right)+(n_{\nu 0}-1)\log\nu-S_{\nu 0}\nu\right\}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\hskip 147.95424pt\left.\left.-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\rho\cdot\exp(h_{t}/2)\cdot\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]^{2}\right\}$
×$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}.$
We approximate the conditional posterior of $\nu$ by a normal distribution centered at its mode and apply a Metropolis-Hastings (MH) step, as done for the generation of $\delta$.
A.1.7 Generation of $\bm{z}_{0}$
The conditional posterior distribution of $z_{0t}$ follows a truncated normal distribution:
$z_{0t}\mid\cdot\sim\mathcal{TN}_{(0,\infty)}(\mu_{zt},\sigma_{zt}^{2}),$
where
$\sigma_{zt}^{-2}=\begin{cases}1+\dfrac{\delta^{2}}{(1-\delta^{2})(1-\rho^{2})},&t=1,\ldots,T-1,\\ \dfrac{1}{1-\delta^{2}},&t=T,\end{cases}$
and
$\mu_{zt}=\begin{cases}\sigma_{zt}^{2}\cdot\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)\delta}{k(\delta)^{2}(1-\rho^{2})\sqrt{1-c^{2}\delta^{2}}}\cdot\left[y_{t}+\sqrt{\dfrac{\lambda_{t}}{\mu_{\lambda}}}\cdot\dfrac{\exp(h_{t}/2)c\delta}{\sqrt{1-c^{2}\delta^{2}}}\right.\\[8.00003pt] \hskip 128.0374pt\left.-\sqrt{\dfrac{\lambda_{t}}{\mu_{\lambda}}}\cdot\rho\cdot\exp(h_{t}/2)\cdot\dfrac{k(\delta)}{\sigma}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right],&t=1,\ldots,T-1,\\[11.99998pt] \sigma_{zT}^{2}\cdot\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{T}}}\cdot\dfrac{\exp(-h_{T}/2)\delta}{k(\delta)^{2}\sqrt{1-c^{2}\delta^{2}}}\cdot\left[y_{T}+\sqrt{\dfrac{\lambda_{T}}{\mu_{\lambda}}}\cdot\dfrac{\exp(h_{T}/2)c\delta}{\sqrt{1-c^{2}\delta^{2}}}\right],&t=T.\end{cases}$
A.1.8 Generation of $\bm{h}$
We first rewrite the RSV-AZ-ST model (12), (13), (14), and (30) as
$x_{t}$ $=\mu_{x}+\alpha_{t}+u_{t},\quad t=1,\ldots,n,$ (67)
$y_{t}$ $=(\delta\bar{z}_{0t}+\sqrt{1-\delta^{2}}z_{t})\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,\quad t=1,\ldots,n,$ (68)
$\alpha_{t+1}$ $=\phi\alpha_{t}+\eta_{t},\quad t=0,\ldots,n-1,$ (69)
where
$\mu_{x}$ $=\xi+\mu,\quad\alpha_{t}=h_{t}-\mu,\quad\bar{z}_{0t}=z_{0t}-c,\quad\gamma=\frac{\exp(\mu/2)}{\sqrt{(1-c^{2}\delta^{2})\mu_{\lambda}}}.$ (70)
The log-likelihood of $(y_{t},x_{t})$ given $\alpha_{t}$, $\alpha_{t+1}$, and other parameters (excluding constants) is
$l_{t}=-\frac{\alpha_{t}}{2}-\frac{(y_{t}-\mu_{t})^{2}}{2\sigma_{t}^{2}}-\frac{(x_{t}-\mu_{x}-\alpha_{t})^{2}}{2\sigma_{u}^{2}},$
where
$\mu_{t}$ =$=\begin{cases}\left[\delta\bar{z}_{0t}+\sqrt{1-\delta^{2}}\dfrac{\rho}{\sigma_{\eta}}(\alpha_{t+1}-\phi\alpha_{t})\right]\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,&t=1,\ldots,n-1,\\ \delta\bar{z}_{0t}\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,&t=n,\end{cases}$ (71)
$\sigma_{t}^{2}$ =$=\begin{cases}(1-\rho^{2})(1-\delta^{2})\lambda_{t}\exp(\alpha_{t})\gamma^{2},&t=1,\ldots,n-1,\\ (1-\delta^{2})\lambda_{t}\exp(\alpha_{t})\gamma^{2},&t=n.\end{cases}$ (72)
Using this log-likelihood, the latent states $(\alpha_{1},\ldots,\alpha_{n})$ can be efficiently sampled via a multi-move sampler based on Omori et al. (2007), Omori and Watanabe (2008, 2015). For further details, see also Takahashi et al. (2016, 2023).
A.1.9 Generation of $\bm{\lambda}$
The conditional probability density function of $\lambda_{t}$ is
$\pi(\lambda_{t}|\cdot)\propto$ $\lambda_{t}^{-(\frac{\nu+1}{2}+1)}$
×$\times\exp\left\{-\frac{\nu}{2\lambda_{t}}-\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
$\left.\left.\hskip 113.81102pt-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\rho\cdot\exp(h_{t}/2)\cdot\frac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\cdot I(t<T)\right]^{2}\right\}.$
We rewrite this density as a product of an inverse gamma kernel and a correction term:
$\pi(\lambda_{t}|\cdot)\propto\lambda_{t}^{-(\frac{a}{2}+1)}\exp\left(-\frac{b_{t}}{2\lambda_{t}}\right)\times g(\lambda_{t}),$
where
$a$ $=\nu+1,$
$b_{t}$ =$=\begin{cases}\nu+\dfrac{\mu_{\lambda}\exp(-h_{t})y_{t}^{2}}{k(\delta)^{2}(1-\rho^{2})},&t=1,\ldots,n-1,\\[8.00003pt] \nu+\dfrac{\mu_{\lambda}\exp(-h_{t})y_{t}^{2}}{k(\delta)^{2}},&t=n,\end{cases}$
$g(\lambda_{t})$ =$=\begin{cases}\exp\left\{\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)y_{t}}{k(\delta)^{2}(1-\rho^{2})}\left[\dfrac{\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}+\rho\cdot\dfrac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]\right\},&t=1,\ldots,n-1,\\[10.00002pt] \exp\left\{\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)y_{t}}{k(\delta)^{2}}\cdot\dfrac{\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right\},&t=n.\end{cases}$
We propose a candidate draw from the inverse gamma distribution,
$\lambda_{t}^{\dagger}\sim\mathcal{IG}\left(\frac{a}{2},\frac{b_{t}}{2}\right),$
and accept it with probability
$\min\left\{1,\frac{g(\lambda_{t}^{\dagger})}{g(\lambda_{t})}\right\}.$
A.2 RSV-FS-ST Model
The joint probability density of $(y_{t},h_{t+1})$ is given by
$f(y_{t},h_{t+1}\mid\bm{\theta},\lambda_{t})=$ $\ q_{T}\left(y_{t}\exp(-h_{t}/2)\mid\gamma,\nu\right)\exp\left(-\frac{1}{2}h_{t}\right)$
×$\times\frac{1}{\sqrt{2\pi(1-\rho^{2})\sigma_{\eta}^{2}}}\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\},$
where $q_{T}$ is defined in (44). Then, the joint posterior density is expressed as
$\pi(\bm{h},\bm{\theta}\mid\bm{x},\bm{y})\propto$ $\ (1-\rho^{2})^{-\frac{T-1}{2}}\sigma_{\eta}^{-T}(1-\phi^{2})^{-\frac{1}{2}}\prod_{t=1}^{T}q_{T}\left(y_{t}\exp(-h_{t}/2)\mid\gamma,\nu\right)\exp\left(-\frac{1}{2}\sum_{t=1}^{T}h_{t}\right)$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
×$\times\exp\left\{-\frac{(1-\phi^{2})}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}\sigma_{u}^{-T}\exp\left\{-\frac{1}{2\sigma_{u}^{2}}\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}\right\}\pi(\bm{\theta}).$
A.2.1 Generation of $\mu$
The conditional posterior distribution of $\mu$ follows a normal distribution:
$\mu\mid\cdot\sim\mathcal{N}(\mu_{1},\sigma_{\mu 1}^{2}),$
where
$\sigma_{\mu 1}^{-2}$ $=\sigma_{\mu 0}^{-2}+\sigma_{\eta}^{-2}\left[1-\phi^{2}+(T-1)(1-\phi)^{2}(1-\rho^{2})^{-1}\right],$
$\mu_{1}$ $=\sigma_{\mu 1}^{2}\left\{\sigma_{\mu 0}^{-2}\mu_{0}+\sigma_{\eta}^{-2}\left[(1-\phi^{2})h_{1}+\frac{1-\phi}{1-\rho^{2}}\sum_{t=1}^{T-1}\left(h_{t+1}-\phi h_{t}-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right)\right]\right\}.$
A.2.2 Generation of $\phi$
The conditional posterior density of $\phi$ is given by
$\pi(\phi\mid\cdot)\propto$ $\ (1-\phi)^{a_{\phi 0}-1}(1+\phi)^{b_{\phi 0}-1}(1-\phi^{2})^{1/2}$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
×$\times\exp\left\{-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
$\propto$ $\ k(\phi)\times\exp\left\{-\frac{1}{2\sigma_{\phi}^{2}}(\phi-\mu_{\phi})^{2}\right\},$
where
$k(\phi)$ $=(1-\phi)^{a_{\phi 0}-1/2}(1+\phi)^{b_{\phi 0}-1/2},$
$\sigma_{\phi}^{2}$ $=\frac{\sigma_{\eta}^{2}(1-\rho^{2})}{\rho^{2}(h_{1}-\mu)^{2}+\sum_{t=2}^{T-1}(h_{t}-\mu)^{2}},$
$\mu_{\phi}$ $=\frac{\sigma_{\phi}^{2}}{\sigma_{\eta}^{2}(1-\rho^{2})}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right](h_{t}-\mu).$
We then employ the MH algorithm, proposing a candidate $\phi^{\dagger}\sim\mathcal{N}(\mu_{\phi},\sigma_{\phi}^{2})$, and accept the candidate with probability $\min\left\{1,k(\phi^{\dagger})/k(\phi)\right\}$.
A.2.3 Generation of $\rho$
The conditional posterior density of $\rho$ is given by
$\pi(\rho\mid\cdot)\propto$ $\ (1+\rho)^{a_{\rho 0}-1}(1-\rho)^{b_{\rho 0}-1}(1-\rho^{2})^{-\frac{T-1}{2}}$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}.$
We approximate the conditional posterior distribution by $\mathcal{N}(\hat{\rho},\sigma_{\hat{\rho}}^{2})$, where $\hat{\rho}$ is the mode of the conditional density and
$\sigma_{\hat{\rho}}^{-2}=-\left.\frac{\partial^{2}\log\pi(\rho\mid\cdot)}{\partial\rho^{2}}\right|_{\rho=\hat{\rho}}.$
We then propose a candidate $\rho^{\dagger}\sim\mathcal{TN}_{(-1,1)}(\hat{\rho},\sigma_{\hat{\rho}}^{2})$, and accept the candidate with probability
$\min\left\{1,\frac{\pi(\rho^{\dagger}\mid\cdot)f_{N}(\rho\mid\hat{\rho},\sigma_{\hat{\rho}}^{2})}{\pi(\rho\mid\cdot)f_{N}(\rho^{\dagger}\mid\hat{\rho},\sigma_{\hat{\rho}}^{2})}\right\}.$
A.2.4 Generation of $\sigma_{\eta}^{2}$
The conditional posterior density of $\sigma_{\eta}^{2}$ is given by
$\pi(\sigma_{\eta}^{2}\mid\cdot)\propto$ $\ (\sigma_{\eta}^{2})^{-\frac{n_{0}+T}{2}-1}\exp\left\{-\frac{S_{0}}{2\sigma_{\eta}^{2}}\right\}\times\exp\left\{-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
$\propto$ $\ k(\sigma_{\eta}^{2})\times(\sigma_{\eta}^{2})^{-\frac{n_{1}}{2}-1}\exp\left\{-\frac{S_{1}}{2\sigma_{\eta}^{2}}\right\},$
where
$n_{1}$ $=n_{0}+T,$
$S_{1}$ $=S_{0}+(1-\phi^{2})(h_{1}-\mu)^{2}+\frac{1}{1-\rho^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)\right]^{2},$
$k(\sigma_{\eta}^{2})$ $=\exp\left\{\frac{\rho}{(1-\rho^{2})\sigma_{\eta}}\sum_{t=1}^{T-1}y_{t}\exp(-h_{t}/2)\left[h_{t+1}-\mu-\phi(h_{t}-\mu)\right]\right\}.$
We then propose a candidate $\sigma_{\eta}^{2\dagger}\sim\mathcal{IG}(n_{1}/2,S_{1}/2)$, and accept the candidate with probability $\min\{1,k(\sigma_{\eta}^{2\dagger})/k(\sigma_{\eta}^{2})\}$.
A.2.5 Generation of $\xi$
The conditional posterior distribution of $\xi$ follows a normal distribution:
$\xi\mid\cdot\sim\mathcal{N}(\mu_{\xi},\sigma_{\xi}^{2}),$
where
$\sigma_{\xi}^{-2}$ $=\sigma_{\xi 0}^{-2}+T\sigma_{u}^{-2},$
$\mu_{\xi}$ $=\sigma_{\xi}^{2}\left[\sigma_{\xi 0}^{-2}\mu_{\xi 0}+\sigma_{u}^{-2}\sum_{t=1}^{T}(x_{t}-h_{t})\right].$
A.2.6 Generation of $\sigma_{u}^{2}$
The conditional posterior distribution of $\sigma_{u}^{2}$ follows an inverse-gamma distribution:
$\sigma_{u}^{2}\mid\cdot\sim\mathcal{IG}\left(\frac{n_{u1}}{2},\frac{S_{u1}}{2}\right),$
where
$n_{u1}$ $=n_{u0}+T,$
$S_{u1}$ $=S_{u0}+\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}.$
A.2.7 Generation of $\gamma$
The conditional posterior density of $\gamma$ is given by
$\pi(\gamma\mid\cdot)\propto$ $\exp\left\{(n_{\gamma 0}-1)\log\gamma-S_{\gamma 0}\gamma+\sum_{t=1}^{T}\log q_{T}\left(y_{t}\exp(-h_{t}/2)\mid\gamma,\nu\right)\right\}.$
We consider the transformation $\tilde{\gamma}=\log\gamma$, and employ a random walk MH algorithm for $\tilde{\gamma}$, where we include a Jacobian adjustment term $\tilde{\gamma}$ in the log conditional posterior density.
A.2.8 Generation of $\nu$
The conditional posterior density of $\nu$ is given by
$\pi(\nu\mid\cdot)\propto\exp\left\{(n_{\nu 0}-1)\log\nu-S_{\nu 0}\nu+\sum_{t=1}^{T}\log q_{T}\left(y_{t}\exp(-h_{t}/2)\mid\gamma,\nu\right)\right\}.$
We consider the transformation $\tilde{\nu}=\log(\nu-2)$, and employ a random walk MH algorithm for $\tilde{\nu}$, where we include a Jacobian adjustment term $\tilde{\nu}$ in the log conditional posterior density.
A.2.9 Generation of $\bm{h}$
First, we define
$g(h_{t}\mid\gamma,\nu)$ $=-\frac{\nu+1}{2}\log\left\{1+\frac{\left(\sigma_{*}y_{t}\exp(-h_{t}/2)+\mu_{*}\right)^{2}}{\nu}\left[\gamma^{-2}I\left(y_{t}\exp(-h_{t}/2)\geq-\frac{\mu_{*}}{\sigma_{*}}\right)\right.\right.$
$\hskip 170.71652pt\left.\left.+\gamma^{2}I\left(y_{t}\exp(-h_{t}/2)<-\frac{\mu_{*}}{\sigma_{*}}\right)\right]\right\}.$
  1. (1)
    For $t=1$, the conditional posterior density of $h_{1}$ is given by
    $\pi(h_{1}\mid\cdot)\propto$ $\ \exp\left\{g(h_{1}\mid\gamma,\nu)-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{2}-\mu-\phi(h_{1}-\mu)-\rho\sigma_{\eta}\exp(-h_{1}/2)y_{1}\right]^{2}\right\}$
    ×$\times\exp\left\{-\frac{1}{2}h_{1}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}-\frac{1}{2\sigma_{u}^{2}}(x_{1}-\xi-h_{1})^{2}\right\}.$
    Define
    $\sigma_{h1}^{-2}$ $=\sigma_{\eta}^{-2}\left(1-\phi^{2}+\frac{\phi^{2}}{1-\rho^{2}}\right)+\sigma_{u}^{-2},$
    $\mu_{h1}$ $=\sigma_{h1}^{2}\left\{-\frac{1}{2}+\frac{\phi\left[h_{2}-(1-\phi)\mu\right]}{\sigma_{\eta}^{2}(1-\rho^{2})}+\frac{(1-\phi^{2})\mu}{\sigma_{\eta}^{2}}+\frac{x_{1}-\xi}{\sigma_{u}^{2}}\right\},$
    $k(h_{1})$ $=\exp\left\{g(h_{1}\mid\gamma,\nu)+\frac{\rho y_{1}\exp(-h_{1}/2)}{(1-\rho^{2})\sigma_{\eta}}\left[h_{2}-\mu-\phi(h_{1}-\mu)\right]-\frac{\rho^{2}y_{1}^{2}\exp(-h_{1})}{2(1-\rho^{2})}\right\}.$
    We then propose a candidate $h_{1}^{\dagger}\sim\mathcal{N}(\mu_{h1},\sigma_{h1}^{2})$, and accept it with probability $\min\{1,k(h_{1}^{\dagger})/k(h_{1})\}$.
  2. (2)
    For $t=2,\ldots,T-1$, the conditional posterior density of $h_{t}$ is given by
    $\pi(h_{t}\mid\cdot)\propto$ $\ \exp\left\{g(h_{t}\mid\gamma,\nu)-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
    ×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t}-\mu-\phi(h_{t-1}-\mu)-\rho\sigma_{\eta}\exp(-h_{t-1}/2)y_{t-1}\right]^{2}\right\}$
    ×$\times\exp\left\{-\frac{1}{2}h_{t}-\frac{1}{2\sigma_{u}^{2}}(x_{t}-\xi-h_{t})^{2}\right\}.$
    Define
    $\sigma_{ht}^{-2}$ $=\sigma_{\eta}^{-2}\left(\frac{1+\phi^{2}}{1-\rho^{2}}\right)+\sigma_{u}^{-2},$
    $\mu_{ht}$ $=\sigma_{ht}^{2}\left\{-\frac{1}{2}+\frac{\phi\left[h_{t+1}-(1-\phi)\mu\right]}{\sigma_{\eta}^{2}(1-\rho^{2})}+\frac{(1-\phi)\mu+\phi h_{t-1}}{\sigma_{\eta}^{2}(1-\rho^{2})}+\frac{\rho y_{t-1}\exp(-h_{t-1}/2)}{(1-\rho^{2})\sigma_{\eta}}+\frac{x_{t}-\xi}{\sigma_{u}^{2}}\right\},$
    $k(h_{t})$ $=\exp\left\{g(h_{t}\mid\gamma,\nu)-\frac{\rho^{2}y_{t}^{2}\exp(-h_{t})}{2(1-\rho^{2})}+\frac{\rho y_{t}\exp(-h_{t}/2)}{(1-\rho^{2})\sigma_{\eta}}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)\right]\right\}.$
    We then propose a candidate $h_{t}^{\dagger}\sim N(\mu_{ht},\sigma_{ht}^{2})$, and accept it with probability $\min\{1,k(h_{t}^{\dagger})/k(h_{t})\}$.
  3. (3)
    For $t=T$, the conditional posterior density of $h_{T}$ is given by
    $\pi(h_{T}\mid\cdot)\propto$ $\ \exp\left\{g(h_{T}\mid\gamma,\nu)-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{T}-\mu-\phi(h_{T-1}-\mu)-\rho\sigma_{\eta}\exp(-h_{T-1}/2)y_{T-1}\right]^{2}\right\}$
    ×$\times\exp\left\{-\frac{1}{2}h_{T}-\frac{1}{2\sigma_{u}^{2}}(x_{T}-\xi-h_{T})^{2}\right\}.$
    Define
    $\sigma_{hT}^{-2}$ $=\sigma_{\eta}^{-2}\left(\frac{1}{1-\rho^{2}}\right)+\sigma_{u}^{-2},$
    $\mu_{hT}$ $=\sigma_{hT}^{2}\left[-\frac{1}{2}+\frac{(1-\phi)\mu+\phi h_{T-1}}{\sigma_{\eta}^{2}(1-\rho^{2})}+\frac{\rho y_{T-1}\exp(-h_{T-1}/2)}{(1-\rho^{2})\sigma_{\eta}}+\frac{x_{T}-\xi}{\sigma_{u}^{2}}\right],$
    $k(h_{T})$ $=\exp\left\{g(h_{T}\mid\gamma,\nu)\right\}.$
    We then propose a candidate $h_{T}^{\dagger}\sim N(\mu_{hT},\sigma_{hT}^{2})$, and accept it with probability $\min\{1,k(h_{T}^{\dagger})/k(h_{T})\}$.